### v1

# Quantifying the Uniqueness of Donald Trump in Presidential Discourse

Abstract
Does Donald Trump speak differently from other presidents? If so, in what ways? Are these differences confined to any single medium of communication? To investigate these questions, this paper introduces a novel metric of uniqueness based on large language models, develops a new lexicon for divisive speech, and presents a framework for comparing the lexical features of political opponents. Applying these tools to a variety of corpora of presidential speeches, we find considerable evidence that Trump’s speech patterns diverge from those of all major party nominees for the presidency in recent history. Some notable findings include Trump’s employment of particularly divisive and antagonistic language targeting of his political opponents and his patterns of repetition for emphasis. Furthermore, Trump is significantly more distinctive than his fellow Republicans, whose uniqueness values are comparably closer to those of the Democrats. These differences hold across a variety of measurement strategies, arise on both the campaign trail and in official presidential addresses, and do not appear to be an artifact of secular time trends.
keywords:
presidential speech, uniqueness metrics, large language models, divisive word lexicon, Donald Trump
1 Introduction
President Donald Trump’s distinctive speech patterns have attracted a considerable amount of media attention.111“The version of English he speaks,” notes Kurt Anderson, “amounts to its own patois, with a special vocabulary and syntax and psychological substrate” [1]. Meanwhile, Andrew Romano makes the point a good deal more plainly: “Donald Trump doesn’t talk like other presidents” [2]. And with good reason. More than just rhetorical flourishes, patterns of presidential speech are constitutive of what Miroff calls the “spectacle” of presidential leadership [3]. Moreover, as both a populist and insurgent within the Republican Party, Trump’s chosen ways of talking are specifically intended to set him apart from other U.S. politicians and enhance his political appeal [4].
A growing body of scholarship examines the characteristics of Trump’s speech patterns, both qualitatively and quantitatively. According to Hart [4], Trump uses language that is explicitly intended to evoke an emotional response from his audience. And other scholars argue that Trump’s speech patterns reveal his autocratic ambitions [5], anti-democratic views [6], a willingness to condone political violence [7], a commitment to populism [8], and low levels of analytic thinking [9]. Quantitative approaches used in these studies are primarily based on lexicons (e.g., [9]) and sentiment classification (e.g., [10]), which necessarily ignore inherent contextual information in speeches. As a result, it is difficult to capture patterns suggested by qualitative studies — e.g., Jamieson and Taussig’s observation that Trump distinguishes himself by being “spontaneous and unpredictable” [6].
Leveraging recent advances in large language models [11], we develop a new quantifier of uniqueness by directly measuring how unpredictable a speech is given the preceding context. In addition, we zoom in on a prevalent yet understudied construct, divisiveness [8]. We operationalize divisiveness as language that is intended to impugn and delegitimize the speaker’s target, and we develop a new lexicon for such speech. Furthermore, we leverage mentions of opponents, especially prevalent in presidential debates, and introduce a comparative framework for the portrayal of political opponents involving lexical features.
We use our proposed tools to analyze a large and diverse corpus of presidential speech. In doing so, we are able to both characterize the overall distinctiveness of Trump’s speech patterns among recent presidential candidates and examine specific qualities that previous research has overlooked. On the campaign trail and in official presidential addresses, we show that Trump is unlike any other recent president. Like previous scholars [12, 13], we show that Trump tends to communicate in shorter, more simplistic sentences. Unlike previous research [9], we find that Trump speaks in ways that are different from all modern presidents, including his immediate predecessors and successor.
Some notable findings from our analysis include his use of distinctive descriptors for his political opponents. Indeed, Trump’s use of divisive and antagonistic language and his tendency to target political opponents distinguish him from all other presidential candidates. Furthermore, Trump stands apart from all his fellow Republicans, whose uniqueness scores are comparatively closer to those of the Democrats. We also document Trump’s patterns of repetition for emphasis.
2 Methodology
2.1 Data
Our research investigates three sources of political speech: presidential debates in general elections since 1960, State of the Union (SOTU) speeches since 1961, and a sample of campaign speeches assembled by the American Presidency Project [14]. Debates and SOTU speeches are generally standard across presidents and time, so all available documents are included in our main analyses. [^2] Additional data were used for model training. For details, see §A.1 and §A.3 for details. Publicly available campaign documents, however, are imbalanced and not comprehensive over the same time frame, so these corpora are limited to speeches delivered within one month of Election Day since 2008. Table 1 provides statistics of the final datasets for which we present results. Of these datasets, Donald Trump spoke in five debates (3,610 sentences), four SOTU addresses (1,471 sentences), and 28 campaign speeches (7,488 sentences) between the years of 2016-2020.
Fig. 1 shows the distribution of sentence lengths from each speaker, across the datasets. Further details of how we collected these datasets can be found in §A.1. From the outset, however, we note that Trump tends to speak in markedly shorter sentences than other presidents do (except Biden). Whereas Trump’s sentences range from 10.4 to 14.5 words in the three data sources, the overall averages ran from 17.6 to 24.4 words. And when comparing presidents within each data source, Trump registered the single lowest number of average words per sentence among all presidents.
Figure 1: Distribution of sentence lengths among all speakers, across debates, SOTU, and campaign data (error bars represent 95%-confidence intervals). Donald Trump tends to use the shortest sentences on average.
DEBATES SOTU CAMPAIGN
# speeches 35 67 187
# sentences 35,096 22,775 36,295
date range 1960-2020 1961-2022 2008-2020
# candidates measured 19 11 6
party ratio (Dem:Rep) 10:9 5:6 3:3
avg. sents/candidate 1,449 2,070 5,701
avg. candidate sent len. 18.76 24.42 17.62
avg. speeches/candidate 3.6 6.1 31.2
Table 1: Overall statistics of the three data types we use to present our main findings.
2.2 Quantifying the uniqueness of political speech
We quantify the uniqueness of political speeches using three complementary approaches: (1) a novel metric based on large language models (LLMs); (2) a new resource of lexicons for divisive speech; and (3) a comparative framework for portrayal of political opponents involving lexical features. Together, these three approaches allow for a robust and multifaceted comparison of speech patterns and shed insights on Trump’s speech patterns both beyond lexical uses and from the perspective of divisiveness. We briefly introduce the intuition behind each approach below.
LLM-based uniqueness
Large language models, like the GPT family of models, have received widespread attention for their abilities to statistically characterize the complex structures of natural language text. LLMs can measure the predictability of text by calculating the likelihood of the next word or sequence of words in a given context, to produce a measure known as perplexity. Perplexity is typically used to evaluate the quality of LLMs [15, 16]. Bits-per-character (BPC) has been used in lieu of perplexity to control for length [17].
Contributing to this literature, we propose a metric of “uniqueness” based on the ability of LLMs to estimate the probabilities of word sequences, and we then use these estimates to compare political speech from various presidential candidates. Specifically, from a pool of candidates, we determine how likely the speech of one speaker is to be produced by the others. A positive value for one candidate’s sentence uniqueness, therefore, suggests that other candidates are unlikely to say it. The advantage of this metric is its consideration of the preceding context of given speech. The precise meaning of the scores, however, may not be interpretable, especially since they only allow for comparisons within a specified pool of speakers. Technical details on the construction of this metric are included in §A.3 and §A.4.
Divisive speech lexicon
To analyze the actual content of language used by candidates, a “divisiveness” lexicon is created and applied to each dataset. We define language as “divisive” if it intends to impugn and delegitimize the speaker’s target, e.g., by attacking their intelligence, integrity, or intentions. Examples of divisive accusations include “racist”, “dishonest”, “corrupt”, or “ridiculous”. Such labels are expressly designed to put the target on defense and accentuate differences and distance between parties. Our definition of divisive is distinct from other commonly analyzed constructs such as political “polarization”, which encompasses language that is associated more with one side than the other but is otherwise agnostic about its valence [18]. Similarly, speech may be divisive without being “toxic,” which contains hateful, abusive, or offensive content [19]. Furthermore, personal attacks can be categorized as a form of toxic language, which prior work has examined primarily in the context of online, written communication [20].
In this study, the divisive word lexicon is used to extract specific examples of divisive language used by presidential candidates. The methodology for creating the lexicon is described in §A.5. The final lexicon consisted of 178 words that four researchers reviewed to be qualitatively “divisive” in the context of political speech.
To the best of our knowledge, we build the first divisive speech lexicon. A strength of this lexicon-based analysis is its easy interpretability and applicability. Divisive words are readily identified and may be used by candidates from any political party in a wide variety of settings. However, the measure is inherently limited by the subjectivity in the creation of the lexicon resource and the lack of contextual consideration, similar to existing lexical approaches.
References to political opponents
Our analysis further expands upon prior work by culling the subset of sentences that explicitly refer to political opponents. We specifically examine presidential debates, in which we define “opponents” here to be either the debate partner or their party. The methodology of tagging opponent mentions is described in §A.2.
Once speech referring to opponents is distinguished, we employ the Fightin’ Words method [21] to identify words more strongly associated with opponent mentions; we extend this comparison across multiple candidates by calculating the overlap of each entity’s corresponding word sets with those of all candidates. Intuitively, a greater overlap indicates that other candidates use similar rhetoric in opponent mentions, while a smaller overlap indicates that other candidates do not use similar rhetoric in opponent mentions. This metric thus provides a novel measure for quantifying a candidate’s distinctiveness with respect to their portrayal of opponents by combining lexical and graph analysis. The results are readily interpretable and clarify the distinctive qualities of language used to characterize one’s opponent. The metric is similarly constrained by the pool of candidates and the finite selection of descriptors.
3 Results
Figure 2: Average sentence uniqueness for each speaker, across all data types. Higher bars indicate greater uniqueness, i.e., that speaker’s speech is less likely to be uttered by other candidates. Trump is the most distinctive speaker among these candidates for debates and SOTU speeches. Error bars correspond to the 95% confidence intervals.
Figure 3: Trump’s average sentence uniqueness scores compared to the others aggregated by party. The number of sentences is denoted on each bar (with error bars representing 95%-confidence intervals). Trump is significantly distinct from other Republicans. The uniqueness scores of the remaining Republicans are closer to those of Democrats than with Trump. For SOTU and campaign speeches, Republicans (without Trump) are less unique than Democrats on average.
3.1 Donald Trump is unique among all presidential candidates in all types of speech
LLM-based uniqueness.
We start by presenting results based on the LLM-based uniqueness metric. We find that Trump is the most distinctive speaker in debates and SOTU speeches (see Figs. 2 and 3). Additionally, among campaign speeches analyzed from 2008 onwards, Trump speaks in ways that stand apart from all other candidates. Some less distinctive speakers include Bill Clinton in debates and George Bush in SOTU; Jimmy Carter is one of the least distinctive speakers for both debates and SOTU. By comparing candidate speech patterns aggregated by party, Fig. 3 shows that Trump is more distinctive than his fellow Republicans for all types of speech. Indeed, the observed difference between Democratic and Republican candidates is minor compared to the gap observed between Trump and everyone else.
While Trump’s speech is characterized by shorter sentences (Fig. 1), we confirm that Trump’s uniqueness is consistent across sentences of all lengths in Fig. 4. We also note that Biden has similarly short sentences on average as Trump, but his uniquness scores are close in magnitude to those of the other candidates, particularly in debates and SOTU.
Furthermore, we find that there is minimal correlation between our uniqueness metric and standard simplicity scores, suggesting that the language model is not conflating uniqueness with language complexity (see §B.2).
Further extensions and robustness checks
When aggregating over presidential terms and years, Trump appears as slightly more distinctive in 2016 than in 2020 for debates, while his SOTU and campaign speeches increase in uniqueness over the years. Overall, he remains consistently more distinctive in both election cycles than the other candidates under consideration (see §B.3). Prior to Trump’s first election cycle, there are no clear temporal trends to suggest that uniqueness has been increasing over time. Moreover, when comparing the uniqueness scores of the top decile of unique sentences for each speaker, we again find that Trump is the most distinctive speaker in all our samples of political speech (see §B.4 ).
Figure 4: Sentence uniqueness across different sentence lengths, for Trump, all other Republicans, and Democrats. For debates and SOTU, Trump is consistently more distinctive across all sentence lengths.
Figure 5: Overall percentage of words used that are divisive. Trump uses the most words from our divisive lexicon, in all types of speech.
3.2 Trump speaks most divisively
Divisive word lexicon
Leveraging our divisive word lexicon, we measure its usage across the different types of presidential speech. The frequency of lexicon word appearances is calculated for each candidate across each dataset and is shown in Fig. 5. Word frequencies are calculated as the number of times a divisive word from the lexicon is spoken by a speaker in a given dataset (debates, SOTU, campaign speeches) divided by the total number of words spoken by the speaker in that in dataset. Both the lexicon words and the speech data are processed to remove contractions and punctuation. Frequency trends provide a macroscopic look at the usage of divisive language over time. In §C, temporal plots and heatmaps of divisive word usage provide a more granular look and do not show strong trends over time.
DEBATES
SOTU
CAMPAIGN
stupid (14),
racist (14),
disgrace (12),
corrupt (8),
disgraceful (8),
ridiculous (6),
ashamed (6),
stupid (6),
filthy (6),
dishonest (6)
cruel (3),
vile (3),
ruthless (2),
foolish (2),
corrupt (2),
reckless (2),
savages (1),
ugly (1),
outrageous (1),
ridiculous (1)
crazy (135),
corrupt (111),
stupid (69),
dishonest (53),
disgrace (45),
ridiculous (31),
racist (27),
incompetent (25),
stupidity (22),
ashamed (22)
Table 2: Top 10 used divisive words by Trump in each type of speech. Raw counts of usage are denoted in parentheses next to each word.
One common trend shown by the frequency plots for each dataset is Donald Trump’s relatively high usage of words from the divisiveness lexicon compared to other speakers. In all three datasets, Donald Trump’s speeches rank highest in divisive word usage compared to the other candidates analyzed. Examples of Trump’s most frequently used words from this lexicon usage are shown in Table 2. He is most verbosely divisive in debates and campaign speeches, uttering terms like “crazy”, “corrupt”, and “stupid” with high frequency. The contexts of debates and campaigns are more combative than SOTU addresses, which is reflected in the higher levels of divisive language in those two mediums. Additional example sentences from various speakers can be found in §C.
DEBATES
Trump: Let me tell you something.
Trump: You take a look at Mosul.
Trump: The biggest problem I have with the stupidity of our foreign policy, we have Mosul.
Trump: I was at a little Haiti the other day in Florida.
Trump: And I want to tell you, they hate the Clintons, because what’s happened in Haiti with the Clinton Foundation is a disgrace.
CAMPAIGN:
Trump: We’re going to bring back the miners and the factory workers and the steel workers.
Trump: We’re going to put them back to work.
Trump: The economic policies of Bill and Hillary Clinton have destroyed manufacturing in your state and throughout the entire country.
Trump: The corrupt Clintons gave us NAFTA.
Trump: The fact is, this is the single most important election in the history of our country.
Trump: And sleepy Joe Biden’s made a corrupt bargain.
Trump: You saw the bargain he made, in exchange for his party’s nomination, which he shouldn’t have gotten because if Pocahontas got out one day early, I’d be running against Crazy Bernie, which would have been okay, too.
Table 3: Example sentences from debates and campaigns, with preceding context, spoken by Trump that use words from the divisive lexicon.
3.3 References to political opponents in speech
In addition to examining overall speech patterns, we can also extract the portions that reference political opponents. Next, we show that Trump is more likely to mention his political opponents than are other presidential candidates and that when doing so, Trump uses particularly distinctive language.
Rates of opponent mentions
As one might expect, candidates routinely mention their opponents in debates and only rarely in SOTU addresses. Overall, rates of sentences that mention opponents for debates, SOTU, and campaign speeches are 20.60%, 0.83%, and 6.95%, respectively (see §D). Trump has the highest rate of opponent mentions in debates.
Since the opponent mentions are more common in presidential debates, we revisit our LLM-based measure for the debates to confirm that Trump is not unique simply because he is more likely to mention opponents. Fig. 6 shows that the LLM-based uniqueness score rankings are relatively consistent across opponent and non-opponent mentions. Moreover, Trump’s language is consistently unique in these debates, whether or not he calls out an opponent.
Figure 6: Uniqueness of speech broken down by opponent mentions, for debates (the error bars represent 95%-confidence intervals). Trump is most distinctive regardless of whether he references an opponent or not.
Fightin’ Words adjective overlap
To identify which adjectives are most strongly associated with describing opponents for each candidate, we calculate the odds ratios between opponent mention sentences and non-opponent mention sentences, i.e., Fightin’ Words (FW) [21]. After these odds ratios are calculated, we compare the words most commonly associated with opponent mentions across candidates and calculate an overlap score. This FW overlap metric for each candidate is calculated as the average number of speakers who share each of the candidate’s top-$n$ FW adjectives. See the Materials and Methods section for further details on the construction of this metric.
[…]
Examples of Trump’s top-$25$ FW adjectives can be seen in Table 4; the left panel includes language used to describe his political opposition, while the right panel includes references to his own party and allies. In addition to attacking opponents (e.g., “disgraceful”), Trump tends to use fairly simplistic adjectives like “massive” and “super”. [^4] Some of these words, like “radical” and “liberal,” appear in the “polarization” dictionary of [18]. Taken as a whole though, our divisive word lexicon contains very little overlap with that dictionary.
(a) FW overlap scores across different top-$n$ thresholds, comparing Trump and the other candidates aggregated by party.
(b) Top-15 FW overlap score in debates for each candidate
Figure 7: Fightin’ Words overlap results. Trump’s FW associated with opponent mentions generally have the lowest overlap in adjective usage compared to other candidates, which is another indication that his language is distinctive.
Mentions Opponents: Y
Mentions Opponents: N
left (2.07), long (1.85), tough (1.80), own (1.79), ok (1.79), bigger (1.79), worse (1.66), radical (1.57), super (1.50), real (1.50), effective (1.47), xenophobic (1.46), liberal (1.46), disgraceful (1.46), massive (1.33), single (1.33), political (1.09), last (1.09), economic (1.08), short (1.08), huge (1.04), various (1.04), red (1.04), upset (1.04), back (1.04)
good (-3.63), great (-3.25), important (-2.36), more (-2.08), inner (-2.00), strong (-1.99), right (-1.84), proud (-1.80), other (-1.79), expensive (-1.64), old (-1.53), big (-1.53), better (-1.50), greatest (-1.44), young (-1.41), sad (-1.41), able (-1.31), much (-1.31), african (-1.27), fine (-1.27), beautiful (-1.27), tougher (-1.18), nice (-1.18), least (-1.12), sure (-1.12)
Table 4: Top-25 FW adjectives said by Donald Trump in debates, either in reference to opponents or not. Z-scores are provided in parentheses.
4 Qualitative analysis of unique political speech
Having found that Trump’s speech is measurably distinct from other politicians, we now further investigate the sources of these differences. Our LLM-based score enables us to identify novel patterns in Trump’s speech. From qualitative analysis of the most and least unique (as determined by our LLM-based score) sentences from Trump, we find that Trump’s speech is distinctive not merely for the particular words that he uses, but for his patterns of repetition and short sentences.
Unique Patterns of Repetition
Here, we focus on a subset of Trump’s speech that reveals a distinctive pattern — his proclivity to repeat a single word for emphasis. In the following examples, a ${}^{\dagger}$ denotes the sentence whose uniqueness score is under assessment. The preceding sentences form the context that the LLM evaluates when scoring the sentence.
The following exchange from a debate between Donald Trump and Hillary Clinton results in one of Trump’s most unique utterances (score 0.23):
Clinton: And when we talk about your business, you’ve taken business bankruptcy six times.
Clinton: There are a lot of great business people that have never taken bankruptcy once.
Clinton: You call yourself the King of Debt.
Clinton: You talk about leverage.
Clinton: You even at one time suggested that you would try to negotiate down the national debt of the United States.
Trump: Wrong.
Trump: ${}^{\dagger}$
This emphatic repetition of “wrong” turns out to be unique to Trump: he is the only candidate to interject “wrong”, which he does 13 times overall in the debates. Furthermore, he specifically repeats the word twice in succession on six separate occasions in his debates, an utterance that is not shared with any other candidate.
This habit of repeating a single word is not confined to the debate setting. One of Trump’s most unique sentences from SOTU speeches include the following from 2019:
Trump: My budget will ask Democrats and Republicans to make the needed commitment to eliminate the HIV epidemic in the United States within 10 years.
Trump: We have made incredible strides.
Trump: ${}^{\dagger}$
Additionally, in 2020:
Trump: I’ve also made an ironclad pledge to American families: We will always protect patients with preexisting conditions.
Trump: And we will always protect your Medicare, and we will always protect your Social Security.
Trump: ${}^{\dagger}$
These examples suggest that a distinguishing feature of Trump’s speech is the use of short phrases that are repeated for emphasis. In fact, Trump has the highest raw count of one-word sentences in all three data types and the highest frequency of one-word sentences in the SOTU and campaign datasets.
5 Discussion
Whether on the campaign trail, the debate stage, or the official dais of the House of Representatives, Trump speaks differently from all modern presidential candidates. We confirm his distinctiveness for speaking in shorter, simpler, and more repetitive sentences. We further quantify the uniqueness of his overall speech compared to that of other presidential candidates. Finally, we demonstrate that Trump uses language that is more divisive, antagonistic, and explicitly focused on his political opponents.
The differences between Trump and other candidates do not appear to be an artifact of secular communication trends (see §B.3), whether in the general coarsening of public discourse or the tendency to speak in simpler sentences. Across multiple points of comparison, observed differences between Trump and other contemporary candidates appear every bit as large as those between Trump and candidates from the 1960s and 70s.
Our findings, of course, come with a variety of limitations. The campaign data, for instance, include samples from only the most recent candidates and exclude certain communication formats, such as interviews. None of our datasets cover instances when surrogates (family members, vice presidents, etc.) speak on behalf of a candidate. While other works have examined Trump’s tweets [22], we do not assess social media postings, due to the lack of comparative data for other candidates, especially older ones. Furthermore, our data collection ends at 2022, excluding more recent remarks from Trump that have been flagged in the media for their escalated divisiveness [23]. While investigating new dimensions of political speech, we recognize that others — e.g., its analytical sophistication or propensity to endorse racial, gender, or sexual stereotypes — warrant still further scrutiny. Also, as previously noted, all of our metrics come with tradeoffs.
Still, the tools we develop can be applied to a wide variety of settings. While we use them to compare speech patterns among presidential candidates, they can just as easily be deployed to analyze the language used by any public officials. And in addition to characterizing overall differences in speech patterns, we illuminate ways of assessing particular qualities that appear in either stand-alone speeches or interactive exchanges.
The substantive findings presented herein, moreover, establish the overwhelming uniqueness of Trump’s speech patterns, just as they reveal particular qualities that distinguish Trump from all modern US presidential candidates. More research is required to map these speech patterns into larger political strategy. Nonetheless, we conjecture that these qualities broadly contribute to Trump’s enduring appeal as a populist who unabashedly denounces established political enemies in a historical period of acute polarization, distrust, and division.
    A.1 Data collection
    Each dataset is scraped from the American Presidents Project (APP) database [14]. In general, all text for each speech is scraped and tagged with its speaker. In addition, some metadata are collected, like the date and title of the speech or document. In order to properly annotate speakers, particularly in the debate speeches, the speaker’s name is identified based on a variety of factors like special text styling, string matching, and html formatting. In addition, audience annotations, like “laughter” and “applause”, are filtered out from the data to the best of our ability. After scraping, the data are subsampled and reviewed to ensure quality and correct speaker annotations.
    Debates data
    The Presidential debates data include General Election debates from 1960 to the present day and Primary Election debates since 2000. For the purposes of this analysis, we exclude primary debates and any general election debates that featured vice-presidential candidates. Sentences spoken by candidates who were not the nominee for either the Republican or Democratic party are also excluded (e.g., Ross Perot in 1992, moderators, and audience members).
    SOTU data
    Because we focus on modern presidents, State of the Union addresses from 1961 onwards are used in deriving our main results. The SOTU dataset includes a few speeches that are not officially considered State of the Union addresses but functionally operate as such. Beginning with Reagan, recent presidents have started addressing a joint session of Congress shortly after their inaugurations. According to the American Presidency Project, “it is probably harmless to categorize these as State of the Union messages (as we do). The impact of such a speech on public, media, and congressional perceptions of presidential leadership and power should be the same as if the address was an official State of the Union”. [^5] https://www.presidency.ucsb.edu/documents/presidential-documents-archive-guidebook/annual-messages-congress-the-state-the-union
    Campaigns data
    All campaign documents from 1932-2020 are collected initially. We manually identify keywords that indicate campaign speeches of interest and filter speech titles based on those keywords. Keywords selected are: “remarks”, “speech”, and “address”. Manually removed documents include: speeches with reporters, question/answer format speeches, press releases, town halls, press releases, co-appearances with other politicians and spouses, etc.
    Subsequently, we apply clique filtering to ensure that any duplicate speeches are removed. It is common that a candidate has a stump speech that is delivered at multiple events, and to diversify and balance the dataset, we remove such duplicates. These de-duplicated data are used for model training.
    Following these filtering steps, our final campaign dataset includes only speeches delivered within the month before election day and only from candidates since 2008. This smaller subset is used to generate the results presented in the main paper.
    A.2 Sentence and opponent mention tagging
    After each dataset is filtered, sentences are tagged as mentioning an opponent or not through an automated process. Sentences are classified as definitely including an opponent mention, possibly including an opponent mention, or not including an opponent mention based on keywords and the presence of parts of speech. For debates, “opponents” are identified by the name(s) of the debate partner or their party. After automatically tagging the debates dataset, the sentences that are labeled as possibly including an opponent mention are manually reviewed by an expert team of four researchers. The team compares a subset of overlapping ratings for consistency. The Cohen’s kappa for inter-coder agreement is roughly 0.8, which indicates substantial agreement.
    For SOTU, opponent mentions only include those of other presidential candidates; names of candidates who never held presidential office are not categorized as mentions of opponents, in contrast to the debates dataset. Following a manual review of selected speeches from the SOTU dataset, it is concluded that references to non-presidential figures are much more neutral and considerably less frequent in State of the Union speeches than they are in debate contexts.
    For campaign data, we automatically tag opponents using the names of the final party candidates from the opposing party of the speaker. Opposing-party candidates from the primaries are not automatically tagged.
    The strict guidelines for automatic tagging ensure high precision of the labels. Limitations of this automatic tagging method include lack of coreference resolution and difficulty of measuring recall.
    A.3 Language model training and analysis
    We fine-tune a pre-trained GPT-2 model using the huggingface [24] and PyTorch Lightning [25] libraries for each of the three data types, resulting in three different model types:
[…]
    To preprocess data for training, we parsed each speech into sentences, prefixed each sentence with the speaker prompt (e.g., “Donald Trump:”, and masked any named entities (as identified by spaCy NER tagger) with a <ENT> mask. [^6] Versions of each model were also trained on unmasked data. The results from these models are consistent with our main findings and are presented in §B.5. We fine-tune each model for 10 epochs on all available corresponding data with a learning rate of $5e$-$5$.
    A.4 BPC/predictability and uniqueness scores
    Consider a set of presidents or candidates $C=\{c_{1},\dots,c_{n}\}$ who each have a corresponding set of sentences $\mathcal{S}=\{S_{1},\dots,S_{n}\}$ where each set $S_{i}=[s_{i}^{1},\dots,s_{i}^{m}]$, for each of the three data types.
    We use bits-per-character (BPC), also known as bits-per-byte, as a proxy for “predictability” of a sentence. Lower BPC values correspond to higher predictability. We use BPC instead of perplexity or loss directly to account for variation in tokenization techniques.
    We calculate BPC as follows. From our fine-tuned models, we are able to obtain loss values, $\mathcal{L}(t_{i})$ for each token $t_{i}$ in an input. For some sentence of tokens $s=(t_{1},t_{2},\dots,t_{k})$ where $\text{len}(s)$ denotes the number of characters in $s$:
    $\mathcal{BPC}(s)=\frac{1}{\text{len}(s)}\sum_{i=1}^{k}\mathcal{L}(t_{i})$
    In other words, $\mathcal{BPC}(s)$ is the sum of cross-entropy losses of each token in a sentence $s$, divided by the number of characters (bytes) in the $s$. We calculate the $\mathcal{BPC}(s)$ of a sentence $s$ with a context window size of 512 tokens. That is, preceding sentences of the one in question are provided to obtain the most representative score of its predictability.
    We first obtain the BPC of a sentence $s_{i}^{j}$ with its original speaker $c_{i}$ as its speaker prompt. For example, “Donald Trump: Make America great again” is denoted by $\mathcal{BPC}_{c_{i}}(s_{i}^{j})$. The BPC of the same sentence $s_{i}^{j}$ with an alternative speaker prompt $c_{k}$, e.g., “Hillary Clinton: Make America great again” is denoted by $\mathcal{BPC}_{c_{k}}(s_{i}^{j})$. Now, we define a “uniqueness” score for $s_{i}^{j}$ sentence as follows:
    $\textsc{SentUniq}(s_{i}^{j})=\left(\frac{1}{|C|-1}\sum_{c_{k}\in C\setminus\{c_{i}\}}\mathcal{BPC}_{c_{k}}(s_{i}^{j})\right)-\mathcal{BPC}_{c_{i}}(s_{i}^{j})$
[…]
    Then, the overall “uniqueness” score for a candidate $c_{i}$ is the average of the SentUniq scores for each of $c_{i}$’s sentences in $S_{i}$:
    $\textsc{Uniq}(c_{i})=\frac{1}{|S_{i}|}\sum_{s_{i}^{j}\in S_{i}}\textsc{SentUniq}(s_{i}^{j})$
    Again, the intuition here is that the larger this score, the greater the average difference is for sentence uniqueness, i.e., overall, speaker $c_{i}$’s sentences are not likely to be said by another candidate.
    A.5 Divisiveness lexicon
    First, a vector space word model is used to populate a list of candidate divisive words. Then, we review and refine the candidate list through researcher annotations. The word vector model used is Gensim’s glove-wiki-gigaword-300 [26, 27].
    Ten seed “divisive” terms are chosen manually by NLP and political science experts as common politically divisive English terms, before seeing the actual speech data. The ten seed terms used are: “stupid”, “dishonest”, “unamerican”, “idiot”, “deplorable”, “pathetic”, “immoral”, “disgrace”, “incompetent”, “foolish”. With these seed terms, 350 additional terms with the highest cosine similarity in the vector space model are added to the lexicon. This initial collection of 360 terms is analyzed by four researchers who annotate each term as “divisive”or “not divisive” based on the following criteria:
    • No modifiers, e.g., “utterly”, “extremely”
    • Should not include words that can often be used both divisively and non-divisively
    • Should only include words that would be considered divisive in most political contexts
    Only words that receive a majority of votes by the annotators are included in the final lexicon. The average pairwise Cohen’s Kappa agreement score of annotations is 0.541, which indicated a moderate level of inter-coder agreement. The final size of the lexicon is 178 words. The entire lexicon can be found in §C.
    Figure 8: Conceptual example of how the FW overlap metric is calculated for N=3.
    A.6 Fightin’ Words overlap metric
    Monroe et al. [21] introduce a methodology for lexical feature selection that calculates odds ratios between of word probabilities between two related corpora. We use this method with the informative Dirichlet prior to obtain the Fightin’ Words in a data type for each candidate.
    For each candidate $s\in C$, we specifically get the set of Fightin’ Words between the words associated with their opponent mentions (${FW}_{Y}(s)$) versus the set of those that are not (${FW}_{N}(s)$). We then examine the top-$n$ of these sets respectively, ${FW}_{Y}^{n}(s)$ and ${FW}_{N}^{n}(s)$. For $Y$ opponent mentions, we create a graph representation as follows (see Fig. 8):
    • speaker nodes $S$, where each node corresponds to a candidate
    • word nodes $W$, where $W$ corresponds to the union of each candidate’s top-$n$ (${FW}_{Y}^{n}(s)$), i.e., $W=\bigcup_{s\in C}{FW}_{Y}^{n}(s)$
    • edges $E(S,W)$, where edge $e$ is added between $s\in S$ and $w\in W$ if $w$ is in $FW_{Y}^{n}(s)$
    The top-$n$ FW overlap metric ($OM_{Y}^{n}(s)$) is then calculated as the
    $OM_{Y}^{n}(s)=\frac{1}{n}\sum_{\forall w\in FW_{Y}^{n}(s)}\operatorname{deg}(w),$
    where $\operatorname{deg}(w)$ corresponds to the degree of the word node (# edges entering $w$). Likewise, for the set of words not associated with opponent mentions, we have
    $OM_{N}^{n}(s)=\frac{1}{n}\sum_{\forall w\in FW_{N}^{n}(s)}\operatorname{deg}(w).$
    Equivalently, the overlap metric for $s$ can be thought of as the average number of speaker sets $FW_{Y}^{n}(c_{i}),~{}\forall c_{i}\in C$ that the top-$n$ words of $s$ appear in. A low $OM_{Y}^{n}(s)$ indicates that speaker $s$ uses more distinct language to refer to opponents, while a higher score corresponds to opponent-referring language that is similar to that of other candidates.
    Appendix B Additional LM-based Uniqueness Results
    (a) Average sentence BPC per candidate (error bars represent 95%-confidence intervals)
    (b) Sentence BPC across sentence length
    Figure 9: Average sentence BPC across each data type. In campaigns, Trump is the most unpredictable among those candidates; however, he is ranks closer to the middle for predictability in debates and SOTU
    (a) Debates
    (b) SOTU
    (c) Campaign
    Figure 10: Heatmaps of Pearson Correlation Coefficients between our uniqueness metric and existing readability scores (FKGL, Gunning Fog, Dale Chall, and ARI). There is little to no correlation between our score and readability; in contrast, there is strong correlation between the different readability indexes.
    B.1 BPC
    Recall that we employ bits-per-character (BPC), also known as bits-per-byte, as a proxy for “predictability” of a sentence. Lower BPC values correspond to higher predictability. We use BPC instead of perplexity or loss directly to account for variation in tokenization techniques. BPC is the sum of cross-entropy losses of each token in a sentence, divided by the number of characters (bytes) in the sentences. Fig. 9 shows average sentence BPC per candidate (Fig. 8(a)) and aggregated by party over sentence length (Fig. 8(b)). While in campaigns, Trump is the most unpredictable among those candidates, he is ranks closer to the middle for predictability in debates and SOTU.
    B.2 Correlation with Readability Metrics
    We calculate the Pearson correlation coefficients for our uniqueness metric with existing readability scores. Our metric has little to no correlation with readability, while different readability indexes show strong correlation with each other (Fig. 10). Fig. 10 also confirms that our metric is dintinct from sentence length.
    B.3 Uniqueness scores over time
    Fig. 11 shows the uniqueness scores over time; in particular, the scores increased in recent years. We also present the uniqueness scores by year/term for debates (Fig. 12), SOTU (Fig. 13), and campaigns (Fig. 14). Overall, Trump is still consistently the most unique every year and term, compared to those of the other presidential candidates. For debates, he is slightly more unique in his first election cycle in 2016, whereas in campaigns, his second election cycle involves more distinctive speech than his first. For SOTU addresses, Trump’s speech becomes more distinct in his second year onwards.
    Figure 11: Uniqueness score over time, aggregated by political party.
    (a) Debates by election year
    (b) Debates by term #
    Figure 12: For debates, Trump is consistently most unique for each term that he runs for president (2016, 2020). He is slightly more unique in his first election cycle, in 2016. The error bars represent 95%-confidence intervals.
    (a) SOTU by term
    (b) SOTU by year
    Figure 13: For SOTU, Trump is consistently most unique throughout all years of his presidency. His speech distinctiveness increases after the first year of his term. The error bars represent 95%-confidence intervals.
    (a) Campaign by election year
    (b) Campaign by term #
    Figure 14: For campaign speeches, Trump is consistently most unique for each term that he runs for president (2016, 2020). His campaign speech in his second election cycle is slightly more unique than his first time. The error bars represent 95%-confidence intervals.
    Figure 15: Uniqueness score by candidate, of sentences within the top 10-ile of scores for each candidate (the error bars represent 95%-confidence intervals). Note that Trump continues to be most unique in all datasets.
    B.4 Top %-ile uniqueness
    Fig. 15 shows that, when comparing the uniqueness scores of the top decile of unique sentences for each speaker, we again find that Trump is the most unique speaker in all our samples of political speech.
    (a) Overall score by candidate
    (b) Scores by sentence length
    (c) Uniqueness by party
    Figure 16: Average sentence uniqueness for each speaker, across all data types, using the UNMASKED model (the error bars represent 95%-confidence intervals). Trump is still the most unique speaker in all types of speeches.
    B.5 LM-based uniqueness results with unmasked model
    In the main paper, we present results using language models that are trained on data with named entities masked out with a <ENT> token, as identified by the spaCy NER tagger. This decision is made to prevent the model from learning patterns like “Only Trump mentions Biden’s name during debates”.
    We also fine-tune each dataset’s model on the data without masking named entities, with results presented in Fig. 16. These results are overall consistent with those of the masked model. Fig. 15(a) shows that Trump remains the most distinctive speaker in all data types, while Fig. 15(b) confirms still that the distinctiveness holds across all sentence lengths. Trump is still more similar in uniqueness to Democrats than his fellow Republicans (Fig. 15(c)).
    Appendix C Additional Divisive Word Lexicon Results
    C.1 Divisive Word Lexicon
    Table 5 contains the 178 words in our proposed divisive word lexicon.
    irresponsible,
    shameful,
    disgraceful,
    disgusting,
    hypocritical,
    idiotic,
    unprofessional,
    ridiculous,
    despicable,
    unethical,
    outrageous,
    arrogant,
    inexcusable,
    ignorant,
    shameless,
    disrespectful,
    absurd,
    ludicrous,
    reprehensible,
    unconscionable,
    scandalous,
    contemptible,
    inept,
    unworthy,
    appalling,
    laughable,
    disingenuous,
    cowardly,
    callous,
    unjust,
    indefensible,
    foolhardy,
    selfish,
    preposterous,
    deceitful
    disloyal,
    ashamed,
    illogical,
    distasteful,
    naive,
    heartless,
    crass,
    repugnant,
    sinful,
    unpatriotic,
    abhorrent,
    obnoxious,
    childish,
    thoughtless,
    stupidity,
    demeaning,
    untrustworthy,
    awful,
    egotistical,
    deluded,
    dumb,
    greedy,
    boorish,
    barbaric,
    tasteless,
    foolishness,
    cruel,
    amoral,
    pretentious,
    hypocrite,
    revolting,
    reckless,
    manipulative,
    nonsensical,
    idiots,
    sexist,
    intolerable,
    uncivilized,
    egregious,
    duplicitous,
    undignified,
    atrocious,
    corrupt,
    contemptuous,
    vile
    liar,
    vindictive,
    careless,
    wretched,
    hateful,
    bigoted,
    bastards,
    undemocratic,
    feckless,
    cowardice,
    pompous,
    mindless,
    spineless,
    arrogance,
    rude,
    treasonous,
    vulgar,
    disgusted,
    insufferable,
    sleazy,
    fool,
    perverse,
    scurrilous,
    insolent,
    grotesque,
    devious,
    incompetence,
    buffoon,
    unintelligent,
    undisciplined,
    odious,
    insane,
    liars,
    depraved,
    crazy,
    moronic,
    uncouth,
    petulant,
    ugly,
    elitist,
    frivolous,
    hypocrisy,
    dastardly,
    incapable,
    tyrannical
    fools,
    heinous,
    thuggish,
    impulsive,
    detestable,
    demented,
    traitor,
    hypocrites,
    filthy,
    conceited,
    traitors,
    loathsome,
    barbarous,
    irrational,
    savages,
    gullible,
    obscene,
    repulsive,
    delusional,
    insult,
    scoundrel,
    abominable,
    deplore,
    mockery,
    perversion,
    bumbling,
    dopey,
    inane,
    nauseating,
    brainless,
    incorrigible,
    exploitive,
    gutless,
    dishonesty,
    unqualified,
    conniving,
    promiscuous,
    degrading,
    racist,
    ruthless,
    perverted,
    diabolical,
    betrayal
    Table 5: The 178 words in our divisive word lexicon, which are verified as being divisive by the majority of 4 annotators. Our original ten seed words are denoted with a $*{}^{*}$.
    C.2 Usage over time
    In debates and campaigns, divisive language usage increased after 2012 (Fig. 17), which corresponds to the onset of Trump’s candidacy. Indeed, we find that Trump uses the most divisive language of all candidates.
    Figure 17: Frequency of usage over time
    (a) Debates
    (b) SOTU
    (c) Campaign
    Figure 18: Heatmaps showing divisive word utterance frequency, normalized over all words spoken, by speaker.
    C.3 Usage by candidate
    Heatmaps of divisive word usage can provide a more granular look at the specific divisive language used by each candidate (Fig. 18). For example, in the debates dataset, it is notable that the frequency of use of the word “racist” has become used more by recent candidates like Trump and Biden. In the debates dataset some of Donald Trump’s most frequently used words in this lexicon are those like “disgrace”, “stupid”, “filthy”, “hate”, and “racist” (Fig. 17(a)). In Donald Trump’s SOTU addresses, the most frequently used divisive words are those like “corrupt”, “vile”, “foolish”, “cruel”, and “savage” (Fig. 17(b)). In campaign speeches, Donald Trump’s most frequently used divisive words are those like “corrupt”, “crazy”, “stupid”, and “dishonest” (Fig. 17(c)).
    C.4 Utterances Using Divisive Words
    Excerpts of speech using divisive words can be found in Table 6, Table 7, and Table 8 for debates, SOTU, and campaign speeches respectively.
    Year Example Sentence(s) w/ Context
    2020
    Joe Biden: I’m talking about the Biden plan. . .
    Donald Trump: Where they want to rip down buildings. . .
    Other: [to Biden] Let him go for a minute, and then you can go.
    Donald Trump: And rebuild the building.
    Joe Biden: No.
    Donald Trump: It’s the dumbest-
    Joe Biden: That is not, that is not. . .
    Donald Trump: … most ridiculous. . .
    2020
    Joe Biden: Russia is paying you a lot.
    Joe Biden: China is paying a lot.
    Joe Biden: And your hotels and all your businesses all around the country, all around the world.
    Joe Biden: And China’s building a new road to a new gas a golf course you have overseas.
    Joe Biden: So what’s going on here?
    Joe Biden: Why don’t you release your tax return or stop talking about corruption?
    2020
    Donald Trump: They were teaching people that our country is a horrible place.
    Donald Trump: It’s a racist place.
    Donald Trump: And they were teaching people to hate our country.
    Donald Trump: And I’m not going to allow that to happen.
    Joe Biden: Nobody’s doing that.
    Other: Vice President Biden.
    Joe Biden: Nobody’s doing that.
    Joe Biden: He’s the racist.
    2020
    Donald Trump: I mean, they can say anything.
    Donald Trump: It’s a very– it makes me sad because I am the least racist person.
    Donald Trump: I can’t even see the audience because it’s so dark, but I don’t care who’s in the audience.
    Donald Trump: I’m the least racist person in this room.
    Other: OK.
    Other: Vice President Biden, let me ask you, very quickly, and then I have a follow up question for you.
    Joe Biden: Abraham Lincoln.
    Joe Biden: Here is one of the most racist presidents we’ve had in modern history.
    Joe Biden: He pours fuel on every single racist fire, every single one.
    2016
    Donald Trump: Well, all of these bad leaders from ISIS are leaving Mosul.
    Donald Trump: Why can’t they do it quietly?
    Donald Trump: Why can’t they do the attack, make it a sneak attack, and after the attack is made, inform the American public that we’ve knocked out the leaders, we’ve had a tremendous success?
    Donald Trump: People leave.
    Donald Trump: Why do they have to say we’re going to be attacking Mosul within the next four to six weeks, which is what they’re saying?
    Donald Trump: How stupid is our country?
    2016
    Donald Trump: But the leaders that we wanted to get are all gone because they’re smart.
    Donald Trump: They say, what do we need this for?
    Donald Trump: So Mosul is going to be a wonderful thing.
    Donald Trump: And Iran should write us a letter of thank you, just like the really stupid the stupidest deal of all time, a deal that’s going to give Iran absolutely nuclear weapons.
    1976
    Gerald Ford: On the other hand, when you have a bill of that magnitude, with those many provisions, a President has to sit and decide if there is more good than bad.
    Gerald Ford: And from the analysis that I have made so far, it seems to me that that tax bill does justify my signature and my approval.
    Other: Governor Carter, your response.
    Jimmy Carter: Well, Mr. Ford is changing considerably his previous philosophy.
    Jimmy Carter: The present tax structure is a disgrace to this country.
    1988
    George Bush: In terms of negative campaigning, you know, I don’t want to sound like a kid in the schoolyard: he started it.
    George Bush: But take a look at the Democratic convention take a look at it.
    George Bush: Do you remember the Senator from Boston chanting out there and the ridicule factor from that lady from Texas that was on there; I mean, come on, this was just outrageous.
    Table 6: [DEBATES] Example usages of words from our divisive lexicon throughout debates.
    Year Example Sentence(s) w/ Context
    2022
    Joe Biden: We’re cutting off Russia’s largest banks from the international financial system; preventing Russia’s Central Bank from defending the Russian ruble, making Putin’s \$630 billion war fund worthless.
    Joe Biden: We’re choking Russia’s access to technology that will sap its economic strength and weaken its military for years to come.
    Joe Biden: Tonight I say to the Russian oligarchs and the corrupt leaders who have bilked billions of dollars off this violent regime: No more.
    2020
    Donald Trump: The terrorist responsible for killing Sergeant Hake was Qasem Soleimani, who provided the deadly roadside bomb that took Chris’s life.
    Donald Trump: Soleimani was the Iranian regime’s most ruthless butcher, a monster who murdered or wounded thousands of American servicemembers in Iraq.
    2019
    Donald Trump: On Friday, it was announced that we added another 304,000 jobs last month alone, almost double the number expected.
    Donald Trump: An economic miracle is taking place in the United States, and the only thing that can stop it are foolish wars, politics, or ridiculous, partisan investigations.
    2018
    Donald Trump: In April, this will be the last time you will ever file under the old and very broken system, and millions of Americans will have more take-home pay starting next month—a lot more.
    Donald Trump: We eliminated an especially cruel tax that fell mostly on Americans making less than \$50,000 a year, forcing them to pay tremendous penalties simply because they couldn’t afford Government-ordered health plans.
    2017
    Donald Trump: As promised, I directed the Department of Defense to develop a plan to demolish and destroy ISIS, a network of lawless savages that have slaughtered Muslims and Christians, and men and women and children of all faiths and all beliefs.
    Donald Trump: We will work with our allies, including our friends and allies in the Muslim world, to extinguish this vile enemy from our planet.
    2016
    Barack Obama: But after years now of record corporate profits, working families won’t get more opportunity or bigger paychecks just by letting big banks or big oil or hedge funds make their own rules at everybody else’s expense.
    Barack Obama: Middle class families are not going to feel more secure because we allowed attacks on collective bargaining to go unanswered.
    Barack Obama: Food stamp recipients did not cause the financial crisis; recklessness on Wall Street did.
    2012
    Barack Obama: Banks had made huge bets and bonuses with other people’s money.
    Barack Obama: Regulators had looked the other way or didn’t have the authority to stop the bad behavior.
    Barack Obama: It was wrong, it was irresponsible, and it plunged our economy into a crisis that put millions out of work, saddled us with more debt, and left innocent, hard-working Americans holding the bag.
    1991
    George Bush: Last year, our friends and allies provided the bulk of the economic costs of Desert Shield.
    George Bush: And now, having received commitments of over \$40 billion for the first 3 months of 1991, I am confident they will do no less as we move through Desert Storm.
    George Bush: But the world has to wonder what the dictator of Iraq is thinking.
    George Bush: If he thinks that by targeting innocent civilians in Israel and Saudi Arabia, that he will gain advantage, he is dead wrong.
    George Bush: If he thinks that he will advance his cause through tragic and despicable environmental terrorism, he is dead wrong.
    1982
    Ronald Reagan: Contrary to some of the wild charges you may have heard, this administration has not and will not turn its back on America’s elderly or America’s poor.
    Ronald Reagan: Under the new budget, funding for social insurance programs will be more than double the amount spent only 6 years ago.
    Ronald Reagan: But it would be foolish to pretend that these or any programs cannot be made more efficient and economical.
    Table 7: [SOTU] Example usages of words from our divisive lexicon throughout SOTU addresses. In general, there are fewer uses of divisive words during SOTU speeches.
    Year Example Sentence(s) w/ Context
    2020
    Joe Biden: The last thing you need is a President who ignores you, looks down at you, who just doesn’t understand you.
    Joe Biden: Like President Trump.
    Joe Biden: His reckless personal conduct since his diagnosis, the destabilizing effect it’s having on our government, is unconscionable.
    2020
    Joe Biden: I’m so grateful to have earned the UA’s endorsement — and to have 355,000 proud plumbers, pipefitters, and more behind me.
    Joe Biden: I also want to thank Rick for sharing his story with us today—and for being part of our convention this year where he shared his story with America.
    Joe Biden: Farmers all across this country have been gutted by President Trump’s broken promises and reckless trade war.
    2020
    Donald Trump: This is the most important election in the history of our country.
    Donald Trump: Six months ago I was saying, ”Well, how do you compare with the last one?”
    Donald Trump: I don’t know.
    Donald Trump: That was important.
    Donald Trump: The fact is, this is the single most important election in the history of our country.
    Donald Trump: And sleepy Joe Biden’s made a corrupt bargain.
    2016
    Donald Trump: The system is rigged.
    Donald Trump: I’ve been saying it for a—it’s rigged, and we’re gonna straighten it out.
    Donald Trump: But the system is rigged.
    Donald Trump: Hillary is not the victim; the American people are the victims of this system.
    Donald Trump: So corrupt in so many ways.
    2016
    Donald Trump: We’re gonna drain the swamp, folks.
    Donald Trump: We’re gonna drain that swamp.
    Donald Trump: Another important issue for Americans is integrity in journalism.
    Donald Trump: These people are among the most dishonest people I’ve ever met, spoken to, done business with.
    Donald Trump: These are the most dishonest people.
    2016
    Donald Trump: We have trade deficits.
    Donald Trump: Think of that.
    Donald Trump: Who negotiates these deals?
    Donald Trump: You know who does?
    Donald Trump: Stupid people.
    Donald Trump: Stupid people.
    Donald Trump: With very stupid leadership.
    2016
    Hillary Clinton: We should honor the men and women in uniform who fight for our country.
    Hillary Clinton: That’s why I was so appalled when Donald Trump tweeted that the new effort underway to push the terrorists out of the key city of Mosul is already, and I quote him, ”a total disaster” and that our country is, again a quote, ”looking dumb.”
    Hillary Clinton: Really?
    Hillary Clinton: He’s declaring defeat before the battle has even started.
    Hillary Clinton: He’s proving once again he is unqualified to be commander in chief of our military.
    2008
    John McCain: In a time of trouble and danger for our country, who will put our country first?
    John McCain: In 21 months, during hundreds of speeches, town halls and debates, I have kept my promise to level with you about my plans to reform Washington and get this country moving again.
    John McCain: As a senator, I’ve seen the corrupt ways of Washington in wasteful spending and other abuses of power.
    2008
    John McCain: In his three short years in the Senate, he has requested nearly a billion dollars in pork projects for his state - a million dollars for every day he’s been in office.
    John McCain: Far from fighting earmarks in Congress, Senator Obama has been an eager participant in this corrupt system.
    Table 8: [CAMPAIGN] Example usages of words from our divisive lexicon throughout campaigns.
    Appendix D Additional Opponent Mention Results
    D.1 Rate of opponent mentions
    The overall rates of sentences that mention opponents for debates, SOTU, and campaign speeches are 20.60%, 0.83%, and 6.95% respectively. Fig. 19 shows the distribution of sentences containing opponent mentions among the candidate speakers. Trump has the highest rate of opponent mentions in debates.
    D.2 Fightin’ Words overlap
    We present additional plots for the Fightin’ Words overlap metric, for different top-$n$ adjectives in debates (see Fig. 7). In particular, for top-$5$, $10$, and $25$ adjective Fightin’ Words, Trump has the lowest overlap in adjectives he uses in association with opponent mentions. These trends are consistent with those presented in the main paper.
    Figure 19: Distribution of sentences containing opponent mentions among all speakers, across the data types. Note the different rates for debates vs. SOTU. Trump has the highest rate of opponent mentions in debates.
    (a) Top-5 FW overlap score in debates for each candidate
    (b) Top-10 FW overlap score in debates for each candidate
    (c) Top-25 FW overlap score in debates for each candidate
    Figure 20: Fightin’ Words overlap results from debates. Trump’s FW associated with opponent mentions generally have the lowest overlap in adjective usage compared to other candidates, which is another indication that his language is distinctive.

### v2

@@
-Quantifying the Uniqueness of Donald Trump in Presidential Discourse
+Quantifying the Uniqueness and Divisiveness of Presidential Discourse
 Abstract
-Does Donald Trump speak differently from other presidents?
+Do American presidents speak discernibly different from each other?
 If so, in what ways?
-Are these differences confined to any single medium of communication?
-To investigate these questions, this paper introduces a novel metric of uniqueness based on large language models, develops a new lexicon for divisive speech, and presents a framework for comparing the lexical features of political opponents.
-Applying these tools to a variety of corpora of presidential speeches, we find considerable evidence that Trump’s speech patterns diverge from those of all major party nominees for the presidency in recent history.
-Some notable findings include Trump’s employment of particularly divisive and antagonistic language targeting of his political opponents and his patterns of repetition for emphasis.
-Furthermore, Trump is significantly more distinctive than his fellow Republicans, whose uniqueness values are comparably closer to those of the Democrats.
-These differences hold across a variety of measurement strategies, arise on both the campaign trail and in official presidential addresses, and do not appear to be an artifact of secular time trends.
+And are these differences confined to any single medium of communication?
+To investigate these questions, this paper introduces a novel metric of uniqueness based on large language models, develops a new lexicon for divisive speech, and presents a framework for assessing the distinctive ways in which presidents speak about their political opponents.
+Applying these tools to a variety of corpora of presidential speeches, we find considerable evidence that Donald Trump’s speech patterns diverge from those of all major party nominees for the presidency in recent history.
+Trump is significantly more distinctive than his fellow Republicans, whose uniqueness values appear closer to those of the Democrats.
+Contributing to these differences is Trump’s employment of divisive and antagonistic language, particularly when targeting his political opponents.
+These differences hold across a variety of measurement strategies, arise on both the campaign trail and in official presidential addresses, and do not appear to be an artifact of secular changes in presidential communications.
 keywords:
 presidential speech, uniqueness metrics, large language models, divisive word lexicon, Donald Trump
+Significance Statement
+While presidential discourse attracts considerable attention from popular media and scholars alike, efforts to computationally compare such rhetoric have been limited to lexical methods.
+This paper proposes a novel suite of metrics to advance such analyses and identify new findings.
+In particular, we leverage large language models to establish an original metric of uniqueness, develop a new lexicon for divisive speech, and introduce a comparative framework for the portrayal of political opponents.
+We then apply these methods to a rich assembly of campaign speeches, presidential debates, and official States of the Union addresses.
+Across all these datasets, we find that Donald Trump’s political rhetoric is unique among modern presidents, and is defined, in part, by his use of antagonistic language, particularly when directed at political opponents.
 1 Introduction
-President Donald Trump’s distinctive speech patterns have attracted a considerable amount of media attention.111“The version of English he speaks,” notes Kurt Anderson, “amounts to its own patois, with a special vocabulary and syntax and psychological substrate” [1].
-Meanwhile, Andrew Romano makes the point a good deal more plainly: “Donald Trump doesn’t talk like other presidents” [2].
-And with good reason.
-More than just rhetorical flourishes, patterns of presidential speech are constitutive of what Miroff calls the “spectacle” of presidential leadership [3].
-Moreover, as both a populist and insurgent within the Republican Party, Trump’s chosen ways of talking are specifically intended to set him apart from other U.S.
-politicians and enhance his political appeal [4].
-A growing body of scholarship examines the characteristics of Trump’s speech patterns, both qualitatively and quantitatively.
-According to Hart [4], Trump uses language that is explicitly intended to evoke an emotional response from his audience.
-And other scholars argue that Trump’s speech patterns reveal his autocratic ambitions [5], anti-democratic views [6], a willingness to condone political violence [7], a commitment to populism [8], and low levels of analytic thinking [9].
-Quantitative approaches used in these studies are primarily based on lexicons (e.g., [9]) and sentiment classification (e.g., [10]), which necessarily ignore inherent contextual information in speeches.
-As a result, it is difficult to capture patterns suggested by qualitative studies — e.g., Jamieson and Taussig’s observation that Trump distinguishes himself by being “spontaneous and unpredictable” [6].
-Leveraging recent advances in large language models [11], we develop a new quantifier of uniqueness by directly measuring how unpredictable a speech is given the preceding context.
-In addition, we zoom in on a prevalent yet understudied construct, divisiveness [8].
+The rise of the modern presidency is defined, in no small measure, by the chief executive’s changing relationship to the American public.
+In what Jeffrey Tulis [1] calls the “rhetorical presidency,” modern presidents are expected to routinely stand before the public in order to explain, persuade, inform, and instruct.
+How precisely they do so, though, is open to interpretation.
+While certain norms of communication govern their behavior, presidents have a fair measure of discretion to speak as they choose.
+When fulfilling their oratory duties, do modern presidents adhere to a common script?
+Or do some presidents defy rhetorical conventions and speak in ways that, at least among themselves, are novel and surprising?
+Leveraging recent advances in large language models (LLMs) [2], we develop a new quantifier of uniqueness by directly measuring the unpredictability of language patterns.
+In addition, we incorporate more standard lexical techniques to examine a prevalent yet understudied construct, divisiveness [3].
 We operationalize divisiveness as language that is intended to impugn and delegitimize the speaker’s target, and we develop a new lexicon for such speech.
-Furthermore, we leverage mentions of opponents, especially prevalent in presidential debates, and introduce a comparative framework for the portrayal of political opponents involving lexical features.
-We use our proposed tools to analyze a large and diverse corpus of presidential speech.
-In doing so, we are able to both characterize the overall distinctiveness of Trump’s speech patterns among recent presidential candidates and examine specific qualities that previous research has overlooked.
-On the campaign trail and in official presidential addresses, we show that Trump is unlike any other recent president.
-Like previous scholars [12, 13], we show that Trump tends to communicate in shorter, more simplistic sentences.
-Unlike previous research [9], we find that Trump speaks in ways that are different from all modern presidents, including his immediate predecessors and successor.
-Some notable findings from our analysis include his use of distinctive descriptors for his political opponents.
-Indeed, Trump’s use of divisive and antagonistic language and his tendency to target political opponents distinguish him from all other presidential candidates.
-Furthermore, Trump stands apart from all his fellow Republicans, whose uniqueness scores are comparatively closer to those of the Democrats.
-We also document Trump’s patterns of repetition for emphasis.
+Furthermore, we introduce a comparative framework for assessing mentions of opponents, which are especially prevalent in presidential debates.
+We use our proposed tools to analyze large and diverse corpora of presidential speech.
+In doing so, we are able to both characterize the overall distinctiveness of presidents’ speech patterns and examine specific qualities that previous research has overlooked.
+Moreover, we can distinguish general speech patterns from those directed towards one’s political opponents.
+In nearly all of our analyses, Donald Trump appears as a clear outlier.
+On the campaign trail, in presidential debates, and in official presidential addresses, we find, Trump’s speech patterns routinely differ from those of all recent presidents — lending credence to Kurt Anderson’s observation that, “The version of English [Trump] speaks amounts to its own patois, with a special vocabulary and syntax and psychological substrate” [4].
+Like previous scholars [5, 6, 7], we find that Trump tends to communicate in shorter, more simplistic sentences.
+But whereas previous quantitative research on presidential rhetoric relied exclusively on lexicons (e.g., [7]) and sentiment classification (e.g., [8]), which necessarily disregard contextual information included in the body of speeches, our methods are able to show that Trump speaks in ways that are holistically different from all modern presidents.
+These differences are pervasive and large — so much so, in fact, that the observed differences between Trump and his fellow Republicans exceed those between Republicans and Democratic presidents.
+Contributing to these differences, we show, is Trump’s tendency to speak in ways that are especially divisive, particularly when focusing on his political opponents.
+[^1] This finding is consistent with previous research showing that Trump uses language that is explicitly intended to evoke an emotional response from his revision audience [9], and that Trump’s speech patterns reveal his autocratic ambitions [10], anti-democratic views [11], a willingness to condone political violence [12], a commitment to populism [3], and low levels of analytic thinking [7].
+None of these scholars, however, have developed a lexicon of divisiveness, as we do in this paper.
+These findings, moreover, are robust to a variety of measurement strategies, arise across rich and diverse corpora of texts, and do not appear to be an artifact of secular time trends.
 2 Methodology
 2.1 Data
-Our research investigates three sources of political speech: presidential debates in general elections since 1960, State of the Union (SOTU) speeches since 1961, and a sample of campaign speeches assembled by the American Presidency Project [14].
+Our research investigates three genres of political speech: presidential debates in general elections since 1960, State of the Union (SOTU) speeches since 1961, and a sample of campaign speeches assembled by the American Presidency Project [13].
 Debates and SOTU speeches are generally standard across presidents and time, so all available documents are included in our main analyses.
-[^2] Additional data were used for model training.
-For details, see §A.1 and §A.3 for details.
-Publicly available campaign documents, however, are imbalanced and not comprehensive over the same time frame, so these corpora are limited to speeches delivered within one month of Election Day since 2008.
-Table 1 provides statistics of the final datasets for which we present results.
-Of these datasets, Donald Trump spoke in five debates (3,610 sentences), four SOTU addresses (1,471 sentences), and 28 campaign speeches (7,488 sentences) between the years of 2016-2020.
+[^2] Additional data are used for model training.
+See §A.3 for details.
+Publicly available campaign documents, however, are imbalanced and not comprehensive over the same time frame, so these corpora are limited to speeches delivered within one month of Election Day in every presidential election since 2008.
+Table 1 provides summary statistics of the final datasets for which we present results.
+[^3] Within these data, Donald Trump spoke in five debates (3,610 sentences), four SOTU addresses (1,471 sentences), and 28 campaign speeches (7,488 sentences) between the years of 2016-2020.
 Fig.
-1 shows the distribution of sentence lengths from each speaker, across the datasets.
-Further details of how we collected these datasets can be found in §A.1.
-From the outset, however, we note that Trump tends to speak in markedly shorter sentences than other presidents do (except Biden).
-Whereas Trump’s sentences range from 10.4 to 14.5 words in the three data sources, the overall averages ran from 17.6 to 24.4 words.
-And when comparing presidents within each data source, Trump registered the single lowest number of average words per sentence among all presidents.
-Figure 1: Distribution of sentence lengths among all speakers, across debates, SOTU, and campaign data (error bars represent 95%-confidence intervals).
-Donald Trump tends to use the shortest sentences on average.
+1 shows the distribution of sentence lengths from each speaker across the datasets.
+Further details of how we collected these datasets can be found in §A.1 From the outset, however, we note that Trump tends to speak in markedly shorter sentences than do other presidents.
+Whereas Trump’s sentences range from 10.4 to 14.5 words in the three data sources, the overall averages range from 17.6 to 24.4 words.
+And when comparing presidents within each data source, Trump registered the single lowest number of average words per sentence among all presidents within debates and campaigns, and the second lowest number within SOTU addresses.
+Figure 1: Distribution of sentence lengths among all speakers, across debates, SOTU, and campaign data (error bars represent 95% CI). Donald Trump tends to use the shortest sentences on average.
+Table 1: Overall statistics of the three data types we use to present our main findings.
 DEBATES SOTU CAMPAIGN
@@
 date range 1960-2020 1961-2022 2008-2020
-# candidates measured 19 11 6
+# candidates compared 19 11 6
 party ratio (Dem:Rep) 10:9 5:6 3:3
@@
 avg. speeches/candidate 3.6 6.1 31.2
-Table 1: Overall statistics of the three data types we use to present our main findings.
 2.2 Quantifying the uniqueness of political speech
-We quantify the uniqueness of political speeches using three complementary approaches: (1) a novel metric based on large language models (LLMs); (2) a new resource of lexicons for divisive speech; and (3) a comparative framework for portrayal of political opponents involving lexical features.
-Together, these three approaches allow for a robust and multifaceted comparison of speech patterns and shed insights on Trump’s speech patterns both beyond lexical uses and from the perspective of divisiveness.
+We quantify the uniqueness of political speeches using three complementary approaches: (i) a novel metric based on LLMs; (ii) a new resource of lexicons for divisive speech; and (iii) a comparative framework for the portrayal of political opponents involving lexical features.
+Together, these three approaches allow for a robust and multifaceted comparison of speech patterns that involve global assessments beyond lexical comparisons and specific evaluations of divisiveness.
 We briefly introduce the intuition behind each approach below.
-LLM-based uniqueness
+LLM-based uniqueness.
 Large language models, like the GPT family of models, have received widespread attention for their abilities to statistically characterize the complex structures of natural language text.
-LLMs can measure the predictability of text by calculating the likelihood of the next word or sequence of words in a given context, to produce a measure known as perplexity.
-Perplexity is typically used to evaluate the quality of LLMs [15, 16].
-Bits-per-character (BPC) has been used in lieu of perplexity to control for length [17].
-Contributing to this literature, we propose a metric of “uniqueness” based on the ability of LLMs to estimate the probabilities of word sequences, and we then use these estimates to compare political speech from various presidential candidates.
-Specifically, from a pool of candidates, we determine how likely the speech of one speaker is to be produced by the others.
+LLMs can measure the predictability of text by calculating the likelihood of the next word or sequence of words in a given context, to produce a measure known as “perplexity,” which is typically used to evaluate the quality of LLMs [14, 15].
+To control for the length of a text, standard perplexity measures can be supplemented with measures of “bits-per-character” (BPC) [16].
+Contributing to this literature, we propose a metric of “uniqueness” based on the ability of LLMs to estimate the probabilities of word sequences, and we then use these estimates to compare political speech from various presidents and presidential candidates, whether delivered from the White House or on the campaign trail.
+Specifically, from a pool of presidential candidates, we determine how likely the speech of one speaker is to be produced by the others.
 A positive value for one candidate’s sentence uniqueness, therefore, suggests that other candidates are unlikely to say it.
+The larger this value, the less likely other candidates are to do so.
 The advantage of this metric is its consideration of the preceding context of given speech.
-The precise meaning of the scores, however, may not be interpretable, especially since they only allow for comparisons within a specified pool of speakers.
-Technical details on the construction of this metric are included in §A.3 and §A.4.
-Divisive speech lexicon
-To analyze the actual content of language used by candidates, a “divisiveness” lexicon is created and applied to each dataset.
+Rather than examine words or phrases in isolation, this metric considers the order in which they appear, and thereby provides a much more nuanced characterization of speech patterns.
+The precise meaning of the scores, however, may be difficult to interpret because they do not reveal the exact features that make a speech distinctive, and because they only allow for comparisons within a specified pool of speakers.
+Technical details on the construction of this metric are included in the §A.3 and §A.4.
+Divisive speech lexicon.
+To analyze the actual content of language used by presidents and presidential candidates, a “divisiveness” lexicon is created and applied to each dataset.
 We define language as “divisive” if it intends to impugn and delegitimize the speaker’s target, e.g., by attacking their intelligence, integrity, or intentions.
@@
 Our definition of divisive is distinct from other commonly analyzed constructs such as political “polarization”, which encompasses language that is associated more with one side than the other but is otherwise agnostic about its valence [17].
-Similarly, speech may be divisive without being “toxic,” which contains hateful, abusive, or offensive content [19].
-Furthermore, personal attacks can be categorized as a form of toxic language, which prior work has examined primarily in the context of online, written communication [20].
-In this study, the divisive word lexicon is used to extract specific examples of divisive language used by presidential candidates.
-The methodology for creating the lexicon is described in §A.5.
-The final lexicon consisted of 178 words that four researchers reviewed to be qualitatively “divisive” in the context of political speech.
-To the best of our knowledge, we build the first divisive speech lexicon.
+Meanwhile, speech may be divisive without being “toxic,” which contains hateful, abusive, or offensive content [18].
+Personal attacks can also be categorized as a form of toxic language, which prior work has examined primarily in the context of online, written communication [19].
+Nor is our measure of divisiveness the simple antonym of traditional notions of politeness [20], which usually involve honoring social conventions, showing gratitude, paying compliments, avoiding complaints and curses, and respecting the listener’s autonomy with the use of softening statements and hedges.
+More than just impolite or insulting, divisive speech, as we conceive it, is explicitly intended to serve the political purposes of delegitimization, marginalization, and distancing between speaker and target.
+To the best of our knowledge, ours is the first divisive speech lexicon.
+As we explain in further detail in §A.5, this lexicon consists of 178 words that four researchers independently reviewed to be qualitatively “divisive” in political speech.
 A strength of this lexicon-based analysis is its easy interpretability and applicability.
-Divisive words are readily identified and may be used by candidates from any political party in a wide variety of settings.
-However, the measure is inherently limited by the subjectivity in the creation of the lexicon resource and the lack of contextual consideration, similar to existing lexical approaches.
-References to political opponents
+Divisive words may be used by candidates from any political party in a wide variety of settings, are readily identified, and, as we subsequently show, are broadly agreed upon by coders.
+As with all lexical approaches, however, the measure is inherently limited by the subjective nature of lexical evaluations and the lack of contextual consideration.
+Our resource, like all lexicons, by itself does not account for surrounding context like negation or valence.
+Strategies like pairing our lexicon with part-of-speech taggers may be leveraged for downstream tasks.
+References to political opponents.
 Our analysis further expands upon prior work by culling the subset of sentences that explicitly refer to political opponents.
@@
 The methodology of tagging opponent mentions is described in §A.2.
-Once speech referring to opponents is distinguished, we employ the Fightin’ Words method [21] to identify words more strongly associated with opponent mentions; we extend this comparison across multiple candidates by calculating the overlap of each entity’s corresponding word sets with those of all candidates.
+Once speech referring to opponents is distinguished, we employ the Fightin’ Words (FW) method [21] to identify words more strongly associated with opponent mentions; we extend this comparison across multiple candidates by calculating the overlap of each entity’s corresponding word sets with those of all presidents and presidential candidates.
 Intuitively, a greater overlap indicates that other candidates use similar rhetoric in opponent mentions, while a smaller overlap indicates that other candidates do not use similar rhetoric in opponent mentions.
@@
 3 Results
-Figure 2: Average sentence uniqueness for each speaker, across all data types.
-Higher bars indicate greater uniqueness, i.e., that speaker’s speech is less likely to be uttered by other candidates.
-Trump is the most distinctive speaker among these candidates for debates and SOTU speeches.
-Error bars correspond to the 95% confidence intervals.
-Figure 3: Trump’s average sentence uniqueness scores compared to the others aggregated by party.
-The number of sentences is denoted on each bar (with error bars representing 95%-confidence intervals).
-Trump is significantly distinct from other Republicans.
-The uniqueness scores of the remaining Republicans are closer to those of Democrats than with Trump.
-For SOTU and campaign speeches, Republicans (without Trump) are less unique than Democrats on average.
 3.1 Donald Trump is unique among all presidential candidates in all types of speech
@@
 Additionally, among campaign speeches analyzed from 2008 onwards, Trump speaks in ways that stand apart from all other candidates.
-Some less distinctive speakers include Bill Clinton in debates and George Bush in SOTU; Jimmy Carter is one of the least distinctive speakers for both debates and SOTU.
 By comparing candidate speech patterns aggregated by party, Fig.
@@
 Indeed, the observed difference between Democratic and Republican candidates is minor compared to the gap observed between Trump and everyone else.
+Figure 2: Average sentence uniqueness for each speaker, across all data types.
+Higher bars indicate greater uniqueness, i.e., that speaker’s speech is less likely to be uttered by other candidates.
+Trump is the most distinctive speaker among these candidates for debates and SOTU speeches.
+Error bars correspond to the 95% CI.
+Figure 3: Trump’s average sentence uniqueness scores compared to the others aggregated by party.
+The number of sentences is denoted on each bar (with error bars representing 95% CI).
+Trump is significantly distinct from other Republicans.
+The uniqueness scores of the remaining Republicans are closer to those of Democrats than with Trump.
+For SOTU and campaign speeches, Republicans (without Trump) are less unique than Democrats on average.
 While Trump’s speech is characterized by shorter sentences (Fig.
@@
 4.
-We also note that Biden has similarly short sentences on average as Trump, but his uniquness scores are close in magnitude to those of the other candidates, particularly in debates and SOTU.
+We also note that Biden has similarly short sentences on average as Trump, but his uniqueness scores are close in magnitude to those of the other candidates, particularly in debates and SOTU addresses.
 Furthermore, we find that there is minimal correlation between our uniqueness metric and standard simplicity scores, suggesting that the language model is not conflating uniqueness with language complexity (see §B.2 ).
-Further extensions and robustness checks
+Further extensions and robustness checks.
 When aggregating over presidential terms and years, Trump appears as slightly more distinctive in 2016 than in 2020 for debates, while his SOTU and campaign speeches increase in uniqueness over the years.
@@
 Figure 5: Overall percentage of words used that are divisive. Trump uses the most words from our divisive lexicon, in all types of speech.
+Score validation.
+The results presented in the main paper are based on the OpenAI GPT-2 model, which was the state-of-the-art open-source causal language model at the start of this project [22].
+We are interested in causal language modeling, i.e., the predictability of text in a forward sequential manner.
+Causal language models are more appropriate for our task of scoring uniqueness, as opposed to masked language models (e.g., BERT, RoBERTa) which are designed to predict missing tokens within a sequence based on bidirectional context (and thus excel at text classification).
+We validate our results with more recent and powerful LLMs like Gemma 2B [23] and Phi1-5b [24] (see §B.6).
+The results from these recent LLMs reaffirm our original findings; that is, Trump is consistently identified as the most unique speaker among modern presidential candidates in all types of speech.
 3.2 Trump speaks most divisively
-Divisive word lexicon
+Divisive word lexicon.
 Leveraging our divisive word lexicon, we measure its usage across the different types of presidential speech.
@@
 In §C, temporal plots and heatmaps of divisive word usage provide a more granular look and do not show strong trends over time.
-DEBATES
-SOTU
-CAMPAIGN
+Table 2: Top 10 divisive words used by Trump in each type of speech. Raw counts of usage are denoted in parentheses next to each word.
+DEBATES SOTU CAMPAIGN
 stupid (14),
@@
 ashamed (22)
-Table 2: Top 10 used divisive words by Trump in each type of speech. Raw counts of usage are denoted in parentheses next to each word.
 One common trend shown by the frequency plots for each dataset is Donald Trump’s relatively high usage of words from the divisiveness lexicon compared to other speakers.
@@
 The contexts of debates and campaigns are more combative than SOTU addresses, which is reflected in the higher levels of divisive language in those two mediums.
-Additional example sentences from various speakers can be found in §C.
-DEBATES
+Table 3 contains a selection of Trump’s divisiveness in context; additional example sentences from various speakers can be found in the §C.
+Across all debates and campaign speeches, which are the two corpora with greatest prevalence of divisive language, sentences that use divisive words are more unique (see §E).
+The Spearman correlation coefficient between uniqueness and divisive word usage is 0.01 in debates and campaigns ($p<0.05$ in campaigns and $p=0.30$ in debates).
+The weak, positive correlation suggests that divisiveness amounts to only a small part of the overall uniqueness of any speech pattern.
+Table 3: Example sentences from debates and campaigns, with preceding context, spoken by Trump that use words from the divisive lexicon. Bolded words are matched with our divisiveness lexicon.
+DEBATES:
 Trump: Let me tell you something.
@@
 Trump: You saw the bargain he made, in exchange for his party’s nomination, which he shouldn’t have gotten because if Pocahontas got out one day early, I’d be running against Crazy Bernie, which would have been okay, too.
-Table 3: Example sentences from debates and campaigns, with preceding context, spoken by Trump that use words from the divisive lexicon.
 3.3 References to political opponents in speech
@@
 Next, we show that Trump is more likely to mention his political opponents than are other presidential candidates and that when doing so, Trump uses particularly distinctive language.
-Rates of opponent mentions
+Rates of opponent mentions.
 As one might expect, candidates routinely mention their opponents in debates and only rarely in SOTU addresses.
@@
 Fig.
-6 shows that the LLM-based uniqueness score rankings are relatively consistent across opponent and non-opponent mentions.
+6 shows that the LLM-based uniqueness score rankings are relatively consistent across opponent and nonopponent mentions.
 Moreover, Trump’s language is consistently unique in these debates, whether or not he calls out an opponent.
-Figure 6: Uniqueness of speech broken down by opponent mentions, for debates (the error bars represent 95%-confidence intervals).
-Trump is most distinctive regardless of whether he references an opponent or not.
-Fightin’ Words adjective overlap
-To identify which adjectives are most strongly associated with describing opponents for each candidate, we calculate the odds ratios between opponent mention sentences and non-opponent mention sentences, i.e., Fightin’ Words (FW) [21].
+Concurrently, sentences that refer to an opponent tend to be more distinct than those that do not (see §E).
+And in debates, we additionally find that sentences that mention opponents contain significantly higher frequencies of divisive words.
+If you are looking to isolate the distinguishing characteristics of a candidate’s speech patterns, therefore, you would do well to focus on how they talk about their political opponents.
+Figure 6: Uniqueness of speech broken down by opponent mentions, for debates (the error bars represent 95% CI). Trump is most distinctive regardless of whether he references an opponent or not.
+Fightin’ Words adjective overlap.
+To identify which adjectives are most strongly associated with describing opponents for each candidate, we calculate the odds ratios between opponent mention sentences and nonopponent mention sentences, i.e., Fightin’ Words (FW) [21].
 After these odds ratios are calculated, we compare the words most commonly associated with opponent mentions across candidates and calculate an overlap score.
 This FW overlap metric for each candidate is calculated as the average number of speakers who share each of the candidate’s top-$n$ FW adjectives.
-See the Materials and Methods section for further details on the construction of this metric.
+See §A.6 for further details on the construction of this metric.
 […]
@@
 Taken as a whole though, our divisive word lexicon contains very little overlap with that dictionary.
-(a) FW overlap scores across different top-$n$ thresholds, comparing Trump and the other candidates aggregated by party.
-(b) Top-15 FW overlap score in debates for each candidate
-Figure 7: Fightin’ Words overlap results.
+Figure 7: FW overlap scores across different top-$n$ thresholds, comparing Trump and the other candidates aggregated by party.
+Figure 8: Top-15 FW overlap score in debates for each candidate.
 Trump’s FW associated with opponent mentions generally have the lowest overlap in adjective usage compared to other candidates, which is another indication that his language is distinctive.
-Mentions Opponents: Y
-Mentions Opponents: N
-left (2.07), long (1.85), tough (1.80), own (1.79), ok (1.79), bigger (1.79), worse (1.66), radical (1.57), super (1.50), real (1.50), effective (1.47), xenophobic (1.46), liberal (1.46), disgraceful (1.46), massive (1.33), single (1.33), political (1.09), last (1.09), economic (1.08), short (1.08), huge (1.04), various (1.04), red (1.04), upset (1.04), back (1.04)
-good (-3.63), great (-3.25), important (-2.36), more (-2.08), inner (-2.00), strong (-1.99), right (-1.84), proud (-1.80), other (-1.79), expensive (-1.64), old (-1.53), big (-1.53), better (-1.50), greatest (-1.44), young (-1.41), sad (-1.41), able (-1.31), much (-1.31), african (-1.27), fine (-1.27), beautiful (-1.27), tougher (-1.18), nice (-1.18), least (-1.12), sure (-1.12)
 Table 4: Top-25 FW adjectives said by Donald Trump in debates, either in reference to opponents or not. Z-scores are provided in parentheses.
-4 Qualitative analysis of unique political speech
-Having found that Trump’s speech is measurably distinct from other politicians, we now further investigate the sources of these differences.
-Our LLM-based score enables us to identify novel patterns in Trump’s speech.
-From qualitative analysis of the most and least unique (as determined by our LLM-based score) sentences from Trump, we find that Trump’s speech is distinctive not merely for the particular words that he uses, but for his patterns of repetition and short sentences.
-Unique Patterns of Repetition
-Here, we focus on a subset of Trump’s speech that reveals a distinctive pattern — his proclivity to repeat a single word for emphasis.
-In the following examples, a ${}^{\dagger}$ denotes the sentence whose uniqueness score is under assessment.
-The preceding sentences form the context that the LLM evaluates when scoring the sentence.
-The following exchange from a debate between Donald Trump and Hillary Clinton results in one of Trump’s most unique utterances (score 0.23):
-Clinton: And when we talk about your business, you’ve taken business bankruptcy six times.
-Clinton: There are a lot of great business people that have never taken bankruptcy once.
-Clinton: You call yourself the King of Debt.
-Clinton: You talk about leverage.
-Clinton: You even at one time suggested that you would try to negotiate down the national debt of the United States.
-Trump: Wrong.
-Trump: ${}^{\dagger}$
-This emphatic repetition of “wrong” turns out to be unique to Trump: he is the only candidate to interject “wrong”, which he does 13 times overall in the debates.
-Furthermore, he specifically repeats the word twice in succession on six separate occasions in his debates, an utterance that is not shared with any other candidate.
-This habit of repeating a single word is not confined to the debate setting. One of Trump’s most unique sentences from SOTU speeches include the following from 2019:
-Trump: My budget will ask Democrats and Republicans to make the needed commitment to eliminate the HIV epidemic in the United States within 10 years.
-Trump: We have made incredible strides.
-Trump: ${}^{\dagger}$
-Additionally, in 2020:
-Trump: I’ve also made an ironclad pledge to American families: We will always protect patients with preexisting conditions.
-Trump: And we will always protect your Medicare, and we will always protect your Social Security.
-Trump: ${}^{\dagger}$
-These examples suggest that a distinguishing feature of Trump’s speech is the use of short phrases that are repeated for emphasis.
-In fact, Trump has the highest raw count of one-word sentences in all three data types and the highest frequency of one-word sentences in the SOTU and campaign datasets.
+Mentions Opponents: Y Mentions Opponents: N
+left (2.07), long (1.85), tough (1.80), own (1.79), ok (1.79), bigger (1.79), worse (1.66), radical (1.57), super (1.50), real (1.50), effective (1.47), xenophobic (1.46), liberal (1.46), disgraceful (1.46), massive (1.33), single (1.33), political (1.09), last (1.09), economic (1.08), short (1.08), huge (1.04), various (1.04), red (1.04), upset (1.04), back (1.04) good (-3.63), great (-3.25), important (-2.36), more (-2.08), inner (-2.00), strong (-1.99), right (-1.84), proud (-1.80), other (-1.79), expensive (-1.64), old (-1.53), big (-1.53), better (-1.50), greatest (-1.44), young (-1.41), sad (-1.41), able (-1.31), much (-1.31), african (-1.27), fine (-1.27), beautiful (-1.27), tougher (-1.18), nice (-1.18), least (-1.12), sure (-1.12)
 4 Discussion
-Whether on the campaign trail, the debate stage, or the official dais of the House of Representatives, Trump speaks differently from all modern presidential candidates.
+In this work, we proposed novel approaches to quantify the uniqueness and divisiveness of presidential discourse.
+Our results show that whether on the campaign trail, the debate stage, or the official dais of the House of Representatives, Trump speaks differently from all modern presidents and presidential candidates.
 We confirm his distinctiveness for speaking in shorter, simpler, and more repetitive sentences.
@@
 The differences between Trump and other candidates do not appear to be an artifact of secular communication trends (see §B.3), whether in the general coarsening of public discourse or the tendency to speak in simpler sentences.
-Across multiple points of comparison, observed differences between Trump and other contemporary candidates appear every bit as large as those between Trump and candidates from the 1960s and 70s.
+Across multiple points of comparison, observed differences between Trump and other contemporary candidates appear every bit as large as those between Trump and candidates from the 1960s and 1970s.
 Our findings, of course, come with a variety of limitations.
@@
 None of our datasets cover instances when surrogates (family members, vice presidents, etc.) speak on behalf of a candidate.
-While other works have examined Trump’s tweets [22], we do not assess social media postings, due to the lack of comparative data for other candidates, especially older ones.
+While other works have examined Trump’s tweets [25], we do not assess any of his social media postings for lack of comparable data for other candidates, especially older ones.
 Furthermore, our data collection ends at 2022, excluding more recent remarks from Trump that have been flagged in the media for their escalated divisiveness [26].
-While investigating new dimensions of political speech, we recognize that others — e.g., its analytical sophistication or propensity to endorse racial, gender, or sexual stereotypes — warrant still further scrutiny.
-Also, as previously noted, all of our metrics come with tradeoffs.
+Our research invites a variety of extensions.
+Given that values of the LLM-based uniqueness metric can only be understood relative to the elements of a selected sample, future work would do well to compare Trump to other speakers, such as populist leaders in other countries.
+Just as we investigate divisiveness as a new dimension of political speech with particular resonance for Trump’s rise to power, other scholars might evaluate other dimensions of political speech, such as its analytical sophistication or propensity to endorse racial, gender, or sexual stereotypes.
+Additionally, future research might explore the possibility of contagion effects from Trump’s speech patterns to other political figures, as politicians periodically emulate and influence one another.
+We also recognize that all of our metrics come with tradeoffs, which we have elucidated throughout this essay.
 Still, the tools we develop can be applied to a wide variety of settings.
-While we use them to compare speech patterns among presidential candidates, they can just as easily be deployed to analyze the language used by any public officials.
+While we use them to compare speech patterns among presidents and presidential candidates, they can just as easily be deployed to analyze the language used by any public officials.
 And in addition to characterizing overall differences in speech patterns, we illuminate ways of assessing particular qualities that appear in either stand-alone speeches or interactive exchanges.
-The substantive findings presented herein, moreover, establish the overwhelming uniqueness of Trump’s speech patterns, just as they reveal particular qualities that distinguish Trump from all modern US presidential candidates.
-More research is required to map these speech patterns into larger political strategy.
+The substantive findings presented herein, moreover, establish the uniqueness of Trump’s speech patterns, just as they reveal particular qualities that distinguish Trump from all modern US presidential candidates.
+More research, of course, is required to map these speech patterns into larger political strategy.
 Nonetheless, we conjecture that these qualities broadly contribute to Trump’s enduring appeal as a populist who unabashedly denounces established political enemies in a historical period of acute polarization, distrust, and division.
+Acknowledgments
+This work is in part supported by a Data & Democracy seed funding at the University of Chicago.
+Data availability
+The data underlying this article are available in the American Presidency Project at http://www.presidency.ucsb.edu/ws, and can be accessed publicly.
+For reproducibility, we release our data and code at https://github.com/ChicagoHAI/quantifying-unique-and-divisive-speech.
+Appendix A Implementation Details
 A.1 Data collection
@@
 Campaigns data
-    All campaign documents from 1932-2020 are collected initially.
+All campaign documents from 1932 to 2020 are collected initially.
 We manually identify keywords that indicate campaign speeches of interest and filter speech titles based on those keywords.
@@
 These de-duplicated data are used for model training.
-    Following these filtering steps, our final campaign dataset includes only speeches delivered within the month before election day and only from candidates since 2008.
+After filtering, our final campaign dataset includes only speeches delivered within the month before election day and only from candidates since 2008.
 This smaller subset is used to generate the results presented in the main paper.
@@
 After automatically tagging the debates dataset, the sentences that are labeled as possibly including an opponent mention are manually reviewed by an expert team of four researchers.
-The team compares a subset of overlapping ratings for consistency.
-The Cohen’s kappa for inter-coder agreement is roughly 0.8, which indicates substantial agreement.
+The team compares a subset of pairwise overlapping ratings for consistency.
+The Cohen’s $\kappa$ for inter-coder agreement is roughly 0.8, which indicates substantial agreement.
 For SOTU, opponent mentions only include those of other presidential candidates; names of candidates who never held presidential office are not categorized as mentions of opponents, in contrast to the debates dataset.
-Following a manual review of selected speeches from the SOTU dataset, it is concluded that references to non-presidential figures are much more neutral and considerably less frequent in State of the Union speeches than they are in debate contexts.
+Following a manual review of selected speeches from the SOTU dataset, it is concluded that references to nonpresidential figures are much more neutral and considerably less frequent in State of the Union speeches than they are in debate contexts.
 For campaign data, we automatically tag opponents using the names of the final party candidates from the opposing party of the speaker.
 Opposing-party candidates from the primaries are not automatically tagged.
-    The strict guidelines for automatic tagging ensure high precision of the labels.
-Limitations of this automatic tagging method include lack of coreference resolution and difficulty of measuring recall.
+The strict guidelines for automatic tagging ensure high precision of the labels. Limitations of this automatic tagging method include lack of coreference resolution and difficulty of measuring recall.
 A.3 Language model training and analysis
-    We fine-tune a pre-trained GPT-2 model using the huggingface [24] and PyTorch Lightning [25] libraries for each of the three data types, resulting in three different model types:
+We fine-tune a pretrained, 124M parameter GPT-2 model using the huggingface [22] and PyTorch Lightning [27] libraries for each of the three data types, resulting in three different language models:
 […]
@@
 [^7] Versions of each model were also trained on unmasked data.
-The results from these models are consistent with our main findings and are presented in §B.5.
-We fine-tune each model for 10 epochs on all available corresponding data with a learning rate of $5e$-$5$.
+The results from these models are consistent with our main findings and are presented in §B.5.) We fine-tune each model for 10 epochs on all available corresponding data with a learning rate of $5e$-$5$.
+Validation with larger, more recent LLMs is shared in §B.6.
 A.4 BPC/predictability and uniqueness scores
@@
 With these seed terms, 350 additional terms with the highest cosine similarity in the vector space model are added to the lexicon.
-This initial collection of 360 terms is analyzed by four researchers who annotate each term as “divisive”or “not divisive” based on the following criteria:
+Each of these initial 360 terms is analyzed by four researchers as “divisive”or “not divisive” based on the following criteria:
   • No modifiers, e.g., “utterly”, “extremely”
-    • Should not include words that can often be used both divisively and non-divisively
+  • Should not include words that can often be used both divisively and nondivisively
   • Should only include words that would be considered divisive in most political contexts
-    Only words that receive a majority of votes by the annotators are included in the final lexicon.
-The average pairwise Cohen’s Kappa agreement score of annotations is 0.541, which indicated a moderate level of inter-coder agreement.
-The final size of the lexicon is 178 words.
-The entire lexicon can be found in §C.
+Only words that receive a majority of votes by the annotators are included in the final lexicon. The final size of the lexicon is 178 words. The entire lexicon can be found in the §C.1.
+Annotator agreement.
+Following [30], in order to analyze how often the annotators agree with each other, we additionally calculate the percentage of times the majority class has size 4 (all annotators agree), size 3 (all but one agree), and size 2 (even split).
+Table 5 shows these agreement values: 58.9% of the 360 initial terms had full annotator agreement, and 86.7% of these 360 have at least 3/4 annotators agreeing on the label.
+When limiting to the 178 terms that at least 3 out of 4 annotators agree are divisive, 69.1% have full agreement on the label (and 100% have majority agreement, by definition).
+So, the final terms include primarily instances that have high agreement scores.
+Table 5: Percentage of majority class annotator agreement on binary labels (divisive?
+“yes” or “no”).
+58.9% of the 360 initial terms had full annotator agreement, and 86.7% of these 360 have at least 3/4 annotators agreeing on the label.
+When limiting to the 178 terms that at least 3/4 annotators agree are divisive, 69.1% have full agreement on the label.
+Majority class size (%)
+Terms four three two
+360 initial 58.9 27.8 13.3
+178 final 69.1 30.9 0.0
+Since there are four annotators, we calculate Fleiss’ $\kappa$ for interannotator agreement, as opposed to Cohen’s $\kappa$ which is designed for two raters.
+For the initial 360 terms, we obtain a Fleiss’ $\kappa$ of $0.54$, which indicates a moderate level of agreement.
+Discussion of this score and why it may underestimate true agreement is included in §C.5.
 Figure 9: Conceptual example of how the FW overlap metric is calculated for $N=3$.
@@
 [21] introduce a methodology for lexical feature selection that calculates odds ratios between of word probabilities between two related corpora.
-We use this method with the informative Dirichlet prior to obtain the Fightin’ Words in a data type for each candidate.
-    For each candidate $s\in C$, we specifically get the set of Fightin’ Words between the words associated with their opponent mentions (${FW}_{Y}(s)$) versus the set of those that are not (${FW}_{N}(s)$).
-We then examine the top-$n$ of these sets respectively, ${FW}_{Y}^{n}(s)$ and ${FW}_{N}^{n}(s)$.
-For $Y$ opponent mentions, we create a graph representation as follows (see Fig.
-8):
+We use this method with the informative Dirichlet prior to obtain the Fightin’ Words (FW) in a data type for each candidate.
+[…]
   • speaker nodes $S$, where each node corresponds to a candidate
-    • word nodes $W$, where $W$ corresponds to the union of each candidate’s top-$n$ (${FW}_{Y}^{n}(s)$), i.e., $W=\bigcup_{s\in C}{FW}_{Y}^{n}(s)$
-    • edges $E(S,W)$, where edge $e$ is added between $s\in S$ and $w\in W$ if $w$ is in $FW_{Y}^{n}(s)$
+[…]
+[…]
 The top-$n$ FW overlap metric ($OM_{Y}^{n}(s)$) is then calculated as the
@@
 $OM_{N}^{n}(s)=\frac{1}{n}\sum_{\forall w\in FW_{N}^{n}(s)}\operatorname{deg}(w).$
-    Equivalently, the overlap metric for $s$ can be thought of as the average number of speaker sets $FW_{Y}^{n}(c_{i}),~{}\forall c_{i}\in C$ that the top-$n$ words of $s$ appear in.
-A low $OM_{Y}^{n}(s)$ indicates that speaker $s$ uses more distinct language to refer to opponents, while a higher score corresponds to opponent-referring language that is similar to that of other candidates.
-    Appendix B Additional LM-based Uniqueness Results
+[…]
+Appendix B LLM-based Uniqueness Robustness Checks
+B.1 BPC
+Recall that we employ bits-per-character (BPC), also known as bits-per-byte, as a measure of “predictability” of a sentence.
+Lower BPC values correspond to higher predictability.
+We use BPC instead of perplexity or loss directly to account for variation in tokenization techniques.
+BPC is the sum of cross-entropy losses of each token in a sentence, divided by the number of characters (bytes) in the sentences.
+Fig.
+10 shows average sentence BPC per candidate (Fig.
+10(a)) and aggregated by party over sentence length (Fig.
+10(b)).
+While in campaigns, Trump is the most unpredictable among those candidates, he is ranks closer to the middle for predictability in debates and SOTU.
+B.2 Correlation with Readability Metrics
+We calculate the Spearman correlation coefficients for our uniqueness metric with existing readability scores.
+Our metric has little to no correlation with readability, while different readability indexes show strong correlation with each other (Fig.
+11).
+Fig.
+11 also confirms that our metric is dintinct from sentence length.
+B.3 Uniqueness scores over time
+Fig.
+12 shows the uniqueness scores over time; in particular, the scores increased in recent years.
+We also present the uniqueness scores by year/term for debates (Fig.
+13), SOTU (Fig.
+14), and campaigns (Fig.
+15).
+Overall, Trump is still consistently the most unique every year and term, compared to those of the other presidential candidates.
+For debates, he is slightly more unique in his first election cycle in 2016, whereas in campaigns, his second election cycle involves more distinctive speech than his first.
+For SOTU addresses, Trump’s speech becomes more distinct in his second year onwards.
+B.4 Top %-ile uniqueness
+Fig.
+16 shows that, when comparing the uniqueness scores of the top decile of unique sentences for each speaker, we again find that Trump is the most unique speaker in all our samples of political speech.
+B.5 LLM-based uniqueness results with unmasked model
+In the main paper, we present results using language models that are trained on data with named entities masked out with a <ENT> token, as identified by the spaCy NER tagger.
+This decision is made to prevent the model from learning patterns like “Only Trump mentions Biden’s name during debates”.
+We also fine-tune each dataset’s model on the data without masking named entities, with results presented in Fig.
+17.
+These results are overall consistent with those of the masked model.
+Fig.
+17(a) shows that Trump remains the most distinctive speaker in all data types, while Fig.
+17(b) confirms still that the distinctiveness holds across all sentence lengths.
+Trump is still more similar in uniqueness to Democrats than his fellow Republicans (Fig.
+17(c)).
+B.6 Validation with other LLMs
+We validate our uniqueness metric implementation with other LLMs.
+Specifically, we recalculate scores with the Gemma-2B and Phi-1.5 (1.3B) base models [23, 24].
+These models are considerably larger than the 124M parameters of GPT-2 [22].
+Like our original setup up for GPT-2, we further train each model on all available corresponding data with a learning rate of $5e$-$5$.
+Each model is trained with LoRA tuning for 15 epochs, using LLaMA-Factory [31].
+As shown in Fig.
+18, while the exact scores differ, the trends are consistent with the main results: that is, Trump’s speech is the most unique among the selection of Democratic and Republican presidential candidates.
+Appendix C Additional Divisive Word Lexicon Results
+C.1 Divisive Word Lexicon
+Table 6 contains the 178 words in our proposed divisive word lexicon.
+C.2 Usage over time
+In debates and campaigns, divisive language usage increased after 2012 (Fig.
+19), which corresponds to the onset of Trump’s candidacy.
+Indeed, we find that Trump uses the most divisive language of all candidates.
+C.3 Usage by candidate
+Heatmaps of divisive word usage can provide a more granular look at the specific divisive language used by each candidate (Fig.
+20).
+For example, in the debates dataset, it is notable that the frequency of use of the word “racist” has become used more by recent candidates like Trump and Biden.
+In the debates dataset some of Donald Trump’s most frequently used words in this lexicon are those like “disgrace”, “stupid”, “filthy”, “hate”, and “racist” (Fig.
+20(a)).
+In Donald Trump’s SOTU addresses, the most frequently used divisive words are those like “corrupt”, “vile”, “foolish”, “cruel”, and “savage” (Fig.
+20(b)).
+In campaign speeches, Donald Trump’s most frequently used divisive words are those like “corrupt”, “crazy”, “stupid”, and “dishonest” (Fig.
+20(c)).
+C.4 Utterances Using Divisive Words
+Excerpts of speech using divisive words can be found in Table 7, Table 8, and Table 9 for debates, SOTU, and campaign speeches respectively.
+C.5 Annotator Agreement
+For the initial 360 terms, we obtain a Fleiss’ $\kappa$ of $0.54$, which indicates a moderate level of agreement.
+Since divisiveness is a somewhat subjective phenomenon, this moderate agreement is similar to that obtained on other subjective lexicons, such as the word-emotion association (average Fleiss’ $\kappa$ of 0.29) [32] and the initial terms in the polarization dictionary (Cohen’s $\kappa$ of 0.61) [17].
+Our final set of 178 terms has full agreement from the majority of annotators, but we still discuss the implications of our Fleiss’ $\kappa$ below.
+Agreement is affected by several characteristics of divisiveness.
+Firstly , divisiveness is a spectrum and terms can be “a little” to “extremely” divisive.
+Annotators only give binary labels, so disagreement may be higher due to lack of nuance.
+Relatedly, the words are presented out of context; they may be more or less divisive depending on usage.
+However, identifying all possible contexts of our lexicon usage is costly and nontrivial.
+Finally, like the word-emotion association lexicon [32], our initial data is biased, with more “not divisive” terms than “divisive” terms (based on only 31.6% of the 360 terms receiving unanimous agreement).
+Both Cohen’s and Fleiss’ $\kappa$ are considered conservative in such settings with label imbalance [32, 33, 34].
+As such, these $\kappa$ values may underestimate the true rater agreement for these cases.
+Ultimately, we introduce the first such lexicon of divisiveness for text analysis.
+Usage of our lexicon may be augmented with additional strategies like dependency parsing and leveraging surrounding context.
+Appendix D Additional Opponent Mention Results
+D.1 Rate of opponent mentions
+The overall rates of sentences that mention opponents for debates, SOTU, and campaign speeches are 20.60%, 0.83%, and 6.95% respectively.
+Fig.
+21 shows the distribution of sentences containing opponent mentions among the candidate speakers.
+Trump has the highest rate of opponent mentions in debates.
+D.2 Fightin’ Words overlap
+We present additional plots for the Fightin’ Words overlap metric, for different top-$n$ adjectives in debates (see Fig.
+22).
+In particular, for top-$5$, $10$, and $25$ adjective Fightin’ Words, Trump has the lowest overlap in adjectives he uses in association with opponent mentions.
+These trends are consistent with those presented in the main paper.
+Appendix E Relating uniqueness, divisiveness, and opponent mentions.
+Fig.
+23(a) shows that in debates and campaign speeches, sentences that use divisive language tend to be more unique as well.
+The Spearman correlation coefficient between divisiveness and uniqueness is 0.01 for debates and campaigns ($p<$0.05 in campaigns and $p=$0.30 in debates).
+For sentences containing opponent mentions in debates, Fig.
+23(b) shows that such utterances tend to be more distinctive and Fig.
+23(c) shows that they tend to have higher divisive word usage as well.
+The Spearman correlation for opponent mentions is 0.05 and 0.02 ($p<$0.05) between uniqueness and divisiveness, respectively.
 (a) Average sentence BPC per candidate (error bars represent 95%-confidence intervals)
@@
 (c) Campaign
-    Figure 10: Heatmaps of Pearson Correlation Coefficients between our uniqueness metric and existing readability scores (FKGL, Gunning Fog, Dale Chall, and ARI).
+Figure 11: Heatmaps of Spearman Correlation Coefficients between our uniqueness metric and existing readability scores (FKGL, Gunning Fog, Dale Chall, and ARI).
 There is little to no correlation between our score and readability; in contrast, there is strong correlation between the different readability indexes.
-    B.1 BPC
-    Recall that we employ bits-per-character (BPC), also known as bits-per-byte, as a proxy for “predictability” of a sentence.
-Lower BPC values correspond to higher predictability.
-We use BPC instead of perplexity or loss directly to account for variation in tokenization techniques.
-BPC is the sum of cross-entropy losses of each token in a sentence, divided by the number of characters (bytes) in the sentences.
-Fig.
-9 shows average sentence BPC per candidate (Fig.
-8(a)) and aggregated by party over sentence length (Fig.
-8(b)).
-While in campaigns, Trump is the most unpredictable among those candidates, he is ranks closer to the middle for predictability in debates and SOTU.
-    B.2 Correlation with Readability Metrics
-    We calculate the Pearson correlation coefficients for our uniqueness metric with existing readability scores.
-Our metric has little to no correlation with readability, while different readability indexes show strong correlation with each other (Fig.
-10).
-Fig.
-10 also confirms that our metric is dintinct from sentence length.
-    B.3 Uniqueness scores over time
-    Fig.
-11 shows the uniqueness scores over time; in particular, the scores increased in recent years.
-We also present the uniqueness scores by year/term for debates (Fig.
-12), SOTU (Fig.
-13), and campaigns (Fig.
-14).
-Overall, Trump is still consistently the most unique every year and term, compared to those of the other presidential candidates.
-For debates, he is slightly more unique in his first election cycle in 2016, whereas in campaigns, his second election cycle involves more distinctive speech than his first.
-For SOTU addresses, Trump’s speech becomes more distinct in his second year onwards.
 Figure 12: Uniqueness score over time, aggregated by political party.
@@
 Note that Trump continues to be most unique in all datasets.
-    B.4 Top %-ile uniqueness
-    Fig.
-15 shows that, when comparing the uniqueness scores of the top decile of unique sentences for each speaker, we again find that Trump is the most unique speaker in all our samples of political speech.
 (a) Overall score by candidate
@@
 Trump is still the most unique speaker in all types of speeches.
-    B.5 LM-based uniqueness results with unmasked model
-    In the main paper, we present results using language models that are trained on data with named entities masked out with a <ENT> token, as identified by the spaCy NER tagger.
-This decision is made to prevent the model from learning patterns like “Only Trump mentions Biden’s name during debates”.
-    We also fine-tune each dataset’s model on the data without masking named entities, with results presented in Fig.
-16.
-These results are overall consistent with those of the masked model.
-Fig.
-15(a) shows that Trump remains the most distinctive speaker in all data types, while Fig.
-15(b) confirms still that the distinctiveness holds across all sentence lengths.
-Trump is still more similar in uniqueness to Democrats than his fellow Republicans (Fig.
-15(c)).
-    Appendix C Additional Divisive Word Lexicon Results
-    C.1 Divisive Word Lexicon
-    Table 5 contains the 178 words in our proposed divisive word lexicon.
+(a) Gemma 2B
+(b) Phi-1.5
+Figure 18: Validation of the LLM-based uniqueness metric with alternative LLMs.
+The resulting trends are consistent with original findings that Trump’s speech is more unique than other Democratic and Republican presidential candidates.
+stupid,
+dishonest,
+unamerican,
+idiot,
+deplorable,
+pathetic,
+immoral,
+disgrace,
+incompetent,
+foolish,
 irresponsible,
@@
 betrayal
-    Table 5: The 178 words in our divisive word lexicon, which are verified as being divisive by the majority of 4 annotators. Our original ten seed words are denoted with a $*{}^{*}$.
-    C.2 Usage over time
-    In debates and campaigns, divisive language usage increased after 2012 (Fig.
-17), which corresponds to the onset of Trump’s candidacy.
-Indeed, we find that Trump uses the most divisive language of all candidates.
+Table 6: The 178 words in our divisive word lexicon, which are verified as being divisive by the majority of 4 annotators. Our original ten seed words are denoted with a .
 Figure 19: Frequency of usage over time
@@
 Figure 20: Heatmaps showing divisive word utterance frequency, normalized over all words spoken, by speaker.
-    C.3 Usage by candidate
-    Heatmaps of divisive word usage can provide a more granular look at the specific divisive language used by each candidate (Fig.
-18).
-For example, in the debates dataset, it is notable that the frequency of use of the word “racist” has become used more by recent candidates like Trump and Biden.
-In the debates dataset some of Donald Trump’s most frequently used words in this lexicon are those like “disgrace”, “stupid”, “filthy”, “hate”, and “racist” (Fig.
-17(a)).
-In Donald Trump’s SOTU addresses, the most frequently used divisive words are those like “corrupt”, “vile”, “foolish”, “cruel”, and “savage” (Fig.
-17(b)).
-In campaign speeches, Donald Trump’s most frequently used divisive words are those like “corrupt”, “crazy”, “stupid”, and “dishonest” (Fig.
-17(c)).
-    C.4 Utterances Using Divisive Words
-    Excerpts of speech using divisive words can be found in Table 6, Table 7, and Table 8 for debates, SOTU, and campaign speeches respectively.
 Year Example Sentence(s) w/ Context
-    2020
-    Joe Biden: I’m talking about the Biden plan. . .
+2020 Joe Biden: I’m talking about the Biden plan. . .
 Donald Trump: Where they want to rip down buildings. . .
@@
 Donald Trump: … most ridiculous. . .
-    2020
-    Joe Biden: Russia is paying you a lot.
+2020 Joe Biden: Russia is paying you a lot.
 Joe Biden: China is paying a lot.
@@
 Joe Biden: Why don’t you release your tax return or stop talking about corruption?
-    2020
+2020 Donald Trump: We have to go back to the core values of this country.
 Donald Trump: They were teaching people that our country is a horrible place.
@@
 Joe Biden: He’s the racist.
-    2020
-    Donald Trump: I mean, they can say anything.
+2020 Donald Trump: I mean, they can say anything.
 Donald Trump: It’s a very– it makes me sad because I am the least racist person.
@@
 Joe Biden: He pours fuel on every single racist fire, every single one.
-    2016
-    Donald Trump: Well, all of these bad leaders from ISIS are leaving Mosul.
+2016 Donald Trump: Well, all of these bad leaders from ISIS are leaving Mosul.
 Donald Trump: Why can’t they do it quietly?
@@
 Donald Trump: How stupid is our country?
-    2016
-    Donald Trump: But the leaders that we wanted to get are all gone because they’re smart.
+2016 Donald Trump: But the leaders that we wanted to get are all gone because they’re smart.
 Donald Trump: They say, what do we need this for?
@@
 Donald Trump: And Iran should write us a letter of thank you, just like the really stupid the stupidest deal of all time, a deal that’s going to give Iran absolutely nuclear weapons.
-    1976
-    Gerald Ford: On the other hand, when you have a bill of that magnitude, with those many provisions, a President has to sit and decide if there is more good than bad.
+2016 Hillary Clinton: And we should demand that Donald release all of his tax returns so that people can see what are the entanglements and the financial relationships that he has…
+Other: We’re going to get to that later.
+Other: Secretary Clinton, you’re out of time.
+Hillary Clinton: … with the Russians and other foreign powers.
+Other: Mr. Trump?
+Donald Trump: Well, I think I should respond, because so ridiculous.
+1976 Gerald Ford: On the other hand, when you have a bill of that magnitude, with those many provisions, a President has to sit and decide if there is more good than bad.
 Gerald Ford: And from the analysis that I have made so far, it seems to me that that tax bill does justify my signature and my approval.
@@
 Jimmy Carter: The present tax structure is a disgrace to this country.
-    1988
-    George Bush: In terms of negative campaigning, you know, I don’t want to sound like a kid in the schoolyard: he started it.
+1988 George Bush: In terms of negative campaigning, you know, I don’t want to sound like a kid in the schoolyard: he started it.
 George Bush: But take a look at the Democratic convention take a look at it.
@@
 Year Example Sentence(s) w/ Context
-    2022
+2022 Joe Biden: Putin is now isolated from the world more than he has ever been.
+Joe Biden: Together, along with our allies, we are right now enforcing powerful economic sanctions.
 Joe Biden: We’re cutting off Russia’s largest banks from the international financial system; preventing Russia’s Central Bank from defending the Russian ruble, making Putin’s \$630 billion war fund worthless.
@@
 Joe Biden: Tonight I say to the Russian oligarchs and the corrupt leaders who have bilked billions of dollars off this violent regime: No more.
-    2020
-    Donald Trump: The terrorist responsible for killing Sergeant Hake was Qasem Soleimani, who provided the deadly roadside bomb that took Chris’s life.
+2020 Donald Trump: The terrorist responsible for killing Sergeant Hake was Qasem Soleimani, who provided the deadly roadside bomb that took Chris’s life.
 Donald Trump: Soleimani was the Iranian regime’s most ruthless butcher, a monster who murdered or wounded thousands of American servicemembers in Iraq.
-    2019
-    Donald Trump: On Friday, it was announced that we added another 304,000 jobs last month alone, almost double the number expected.
+2019 Donald Trump: On Friday, it was announced that we added another 304,000 jobs last month alone, almost double the number expected.
 Donald Trump: An economic miracle is taking place in the United States, and the only thing that can stop it are foolish wars, politics, or ridiculous, partisan investigations.
-    2018
-    Donald Trump: In April, this will be the last time you will ever file under the old and very broken system, and millions of Americans will have more take-home pay starting next month—a lot more.
+2018 Donald Trump: I am asking Congress to address the fundamental flaws in the terrible Iran nuclear deal.
+Donald Trump: My administration has also imposed tough sanctions on the communist and socialist dictatorships in Cuba and Venezuela.
+Donald Trump: But no regime has oppressed its own citizens more totally or brutally than the cruel dictatorship in North Korea.
+2018 Donald Trump: In April, this will be the last time you will ever file under the old and very broken system, and millions of Americans will have more take-home pay starting next month—a lot more.
 Donald Trump: We eliminated an especially cruel tax that fell mostly on Americans making less than \$50,000 a year, forcing them to pay tremendous penalties simply because they couldn’t afford Government-ordered health plans.
-    2017
+2017 Donald Trump: We cannot allow a beachhead of terrorism to form inside America.
+Donald Trump: We cannot allow our Nation to become a sanctuary for extremists.
+Donald Trump: That is why my administration has been working on improved vetting procedures, and we will shortly take new steps to keep our Nation safe and to keep those out who will do us harm.
 Donald Trump: As promised, I directed the Department of Defense to develop a plan to demolish and destroy ISIS, a network of lawless savages that have slaughtered Muslims and Christians, and men and women and children of all faiths and all beliefs.
 Donald Trump: We will work with our allies, including our friends and allies in the Muslim world, to extinguish this vile enemy from our planet.
-    2016
-    Barack Obama: But after years now of record corporate profits, working families won’t get more opportunity or bigger paychecks just by letting big banks or big oil or hedge funds make their own rules at everybody else’s expense.
+2016 Barack Obama: But after years now of record corporate profits, working families won’t get more opportunity or bigger paychecks just by letting big banks or big oil or hedge funds make their own rules at everybody else’s expense.
 Barack Obama: Middle class families are not going to feel more secure because we allowed attacks on collective bargaining to go unanswered.
 Barack Obama: Food stamp recipients did not cause the financial crisis; recklessness on Wall Street did.
-    2012
+2012 Barack Obama: In 2008, the house of cards collapsed.
+Barack Obama: We learned that mortgages had been sold to people who couldn’t afford or understand them.
 Barack Obama: Banks had made huge bets and bonuses with other people’s money.
@@
 Barack Obama: It was wrong, it was irresponsible, and it plunged our economy into a crisis that put millions out of work, saddled us with more debt, and left innocent, hard-working Americans holding the bag.
-    1991
-    George Bush: Last year, our friends and allies provided the bulk of the economic costs of Desert Shield.
+1991 George Bush: Last year, our friends and allies provided the bulk of the economic costs of Desert Shield.
 George Bush: And now, having received commitments of over \$40 billion for the first 3 months of 1991, I am confident they will do no less as we move through Desert Storm.
@@
 George Bush: If he thinks that he will advance his cause through tragic and despicable environmental terrorism, he is dead wrong.
-    1982
-    Ronald Reagan: Contrary to some of the wild charges you may have heard, this administration has not and will not turn its back on America’s elderly or America’s poor.
+1982 Ronald Reagan: Contrary to some of the wild charges you may have heard, this administration has not and will not turn its back on America’s elderly or America’s poor.
 Ronald Reagan: Under the new budget, funding for social insurance programs will be more than double the amount spent only 6 years ago.
@@
 Year Example Sentence(s) w/ Context
-    2020
-    Joe Biden: The last thing you need is a President who ignores you, looks down at you, who just doesn’t understand you.
+2020 Joe Biden: The last thing you need is a President who ignores you, looks down at you, who just doesn’t understand you.
 Joe Biden: Like President Trump.
 Joe Biden: His reckless personal conduct since his diagnosis, the destabilizing effect it’s having on our government, is unconscionable.
-    2020
-    Joe Biden: I’m so grateful to have earned the UA’s endorsement — and to have 355,000 proud plumbers, pipefitters, and more behind me.
+2020 Joe Biden: I’m so grateful to have earned the UA’s endorsement — and to have 355,000 proud plumbers, pipefitters, and more behind me.
 Joe Biden: I also want to thank Rick for sharing his story with us today—and for being part of our convention this year where he shared his story with America.
 Joe Biden: Farmers all across this country have been gutted by President Trump’s broken promises and reckless trade war.
-    2020
-    Donald Trump: This is the most important election in the history of our country.
+2020 Donald Trump: This is the most important election in the history of our country.
 Donald Trump: Six months ago I was saying, ”Well, how do you compare with the last one?”
@@
 Donald Trump: And sleepy Joe Biden’s made a corrupt bargain.
-    2016
+2016 Donald Trump: Now, Bernie Sanders should be angry right?
+Donald Trump: Shouldn’t he be angry?
+Donald Trump: Now, I’ll tell you what.
+Donald Trump: The system is rigged.
 Donald Trump: The system is rigged.
@@
 Donald Trump: So corrupt in so many ways.
-    2016
+2016 Donald Trump: I want the entire corrupt Washington establishment to hear and to heed the words we all will be saying right now.
+Donald Trump: When we win on November 8th, we are going to Washington, D.C. and we are going to drain the swamp.
+Donald Trump: Gonna drain the swamp.
 Donald Trump: We’re gonna drain the swamp, folks.
@@
 Donald Trump: These are the most dishonest people.
-    2016
+2016 Donald Trump: Our trade deals, we lose \$800 billion a year on trade.
 Donald Trump: We have trade deficits.
@@
 Donald Trump: With very stupid leadership.
-    2016
-    Hillary Clinton: We should honor the men and women in uniform who fight for our country.
+2016 Donald Trump: We will terminate NAFTA and get a much better deal for our workers if we can’t renegotiate it properly.
+Donald Trump: We’re gonna get a better deal for our workers and for our companies.
+Donald Trump: Because we cannot continue to be the people led by stupid people.
+2016 Hillary Clinton: We should honor the men and women in uniform who fight for our country.
 Hillary Clinton: That’s why I was so appalled when Donald Trump tweeted that the new effort underway to push the terrorists out of the key city of Mosul is already, and I quote him, ”a total disaster” and that our country is, again a quote, ”looking dumb.”
@@
 Hillary Clinton: He’s proving once again he is unqualified to be commander in chief of our military.
-    2008
-    John McCain: In a time of trouble and danger for our country, who will put our country first?
+2016 Hillary Clinton: But this is not new.
+Hillary Clinton: I know I’m reaching out to Republicans and Independents as well as Democrats because I want to be the president for all Americans.
+Hillary Clinton: And, when you think about it, what he said at the convention, ’I alone can fix it,’ runs counter to who we are as Americans.
+Hillary Clinton: We work together.
+Hillary Clinton: So there are many reasons why I think it is fair to conclude that Donald Trump is unqualified and unfit to be president.
+2008 John McCain: In a time of trouble and danger for our country, who will put our country first?
 John McCain: In 21 months, during hundreds of speeches, town halls and debates, I have kept my promise to level with you about my plans to reform Washington and get this country moving again.
 John McCain: As a senator, I’ve seen the corrupt ways of Washington in wasteful spending and other abuses of power.
-    2008
-    John McCain: In his three short years in the Senate, he has requested nearly a billion dollars in pork projects for his state - a million dollars for every day he’s been in office.
+2008 John McCain: In his three short years in the Senate, he has requested nearly a billion dollars in pork projects for his state - a million dollars for every day he’s been in office.
 John McCain: Far from fighting earmarks in Congress, Senator Obama has been an eager participant in this corrupt system.
 Table 9: [CAMPAIGN] Example usages of words from our divisive lexicon throughout campaigns.
-    Appendix D Additional Opponent Mention Results
-    D.1 Rate of opponent mentions
-    The overall rates of sentences that mention opponents for debates, SOTU, and campaign speeches are 20.60%, 0.83%, and 6.95% respectively.
-Fig.
-19 shows the distribution of sentences containing opponent mentions among the candidate speakers.
-Trump has the highest rate of opponent mentions in debates.
-    D.2 Fightin’ Words overlap
-    We present additional plots for the Fightin’ Words overlap metric, for different top-$n$ adjectives in debates (see Fig.
-7).
-In particular, for top-$5$, $10$, and $25$ adjective Fightin’ Words, Trump has the lowest overlap in adjectives he uses in association with opponent mentions.
-These trends are consistent with those presented in the main paper.
 Figure 21: Distribution of sentences containing opponent mentions among all speakers, across the data types.
@@
 Trump’s FW associated with opponent mentions generally have the lowest overlap in adjective usage compared to other candidates, which is another indication that his language is distinctive.
+(a) Sentence uniqueness across divisive language usage, for each type of speech.
+(b) Sentence uniqueness across opponent mentions, in debates.
+(c) Divisive language frequency across opponent mentions, in debates
+Figure 23: Pairwise analysis of our three proposed metrics, aggregated globally across all speakers.
+Fig.
+23(a) shows that in debates and campaign speeches, sentences that use divisive language tend to be more unique as well.
+For sentences containing opponent mentions in debates, Fig.
+23(b) shows that such utterances tend to be more distinctive and Fig.
+23(c) shows that they tend to have higher divisive word usage as well.